### v1

# On the mixed UDA states and additivity

Abstract
Mixed states that are uniquely determined among all (UDA) states are vital in efficient quantum tomography. We show the necessary and sufficient conditions by which some multipartite mixed states are UDA by their $k$-partite reduced density matrices. The case for $k=2$ is mostly studied, which requires minimal local information and shows practical benefits. Based on that, we establish a systematic method for determining UDA states and provide a complete characterization of the additivity of UDA bipartite and three-qubit product states. We show the application of mixed UDA states and their characterization from the perspectives of tomography and other tasks.
I Introduction
Understanding the information of a global state with its local parts elucidates profound differences between quantum mechanics and its classical counterpart. Quantum marginal problem deals with the existence of a valid global state, whose marginals (i.e., reduced density matrices) match the input set of states coleman1963structure; klyachko2004quantum. If such a quantum marginal problem does admit a solution, one may focus on an interrelated problem. Given access to local information provided by the marginals, is the global state determined uniquely by the marginals? If yes, this global state is uniquely determined among all (UDA) states, which naturally follows the need of quantum system certification, including quantum state tomography (QST) bisio2009optimal; huang2025process, self-testing ivan2023quantum, and other device-independent tasks qiu2024measurement. Generally, the number of observables required for certification increases exponentially with the number of systems qin2024state, whereas the UDA property allows the required observables exhibits polynomial growth. This lowers the complexity of reconstructing the state of systems substantially cramer2010efficient; experiment2024hu. Besides, the UDA property brings tools for dissipative quantum control karuvade2019uniquely, and topological stabilizer codes in multipartite quantum error correction liao2022topo. Moreover, the characterization of a set of marginals that uniquely determines the global state sheds light on the structure of multipartite entanglement across different subsystems den2002almost; walck2008only.
The UDA problem arises from the scope of pure states. Given a pure state $|\psi\rangle$, if there are no other states sharing the same marginals with $|\psi\rangle$, it is said to be uniquely determined. This is further clarified into two categories: uniquely determined among pure (UDP) states and uniquely determined among all (UDA) states. By the two categories, pure states can be neither UDP nor UDA, UDP but not UDA, or UDA states qt2017tao. The characterization of pure, uniquely determined states has been extensively studied. Almost all three-qubit pure states, general symmetric Dicke and $W$-type states have been proved to be UDA by 2-marginals den2002almost; rana2011optimal; yu2013multi; zhang2024EvenParticle. The only $n$-particle pure states which can not be UDP by their (n1)(n-1)-marginals are those GHZ-type states, and the result is further improved to the case of UDA walck2008only. In addition to the pure states, mixed states can also be uniquely determined among all states, for example, a tripartite UDA state with equal-dimensional subsystems has been constructed role2014chen. Mixed states exist commonly in laboratories than pure ones since real quantum systems inevitably interact with their environment, and such decoherence and noise destroy purity liu2023observation; lee2025topo. From this perspective, the characterization of mixed UDA states has both fundamental and practical implications. Compared with pure states, UDA mixed states have been less explored. The certification of mixed UDA states is not easy, even when the state is given numerically. It is shown that almost every tripartite density operator with low rank is UDA by two specific bipartite marginals, and thus almost all four-qubit states of rank two are UDA by tripartite marginals uinque2013chen. For the convenience of simplification, the scope of some existing results is focused on ”almost all states”, by excluding a measure-zero set of states jones2005parts; wyder2017almost; parashar2009nqubit. However, for a state under consideration, it is not straightforward to verify whether it is in the measure-zero set. Therefore, it is of practical interest to give a complete characterization of mixed UDA states by giving the necessary and sufficient conditions. This is the first motivation of this work.
For a set of marginals on subsystems of fixed size, the global state may transition from a UDA state to a non-UDA state as the number of systems increases. For example, generic states of $n$-qudits are UDA by certain sets of marginals of less than about two-thirds linden2002parts, or more than half of the parties jones2005parts, whereas the reduced states of fewer than half of the parties are not sufficient. This implies that given two mixed UDA states $\alpha$ and $\beta$, their additive form $\alpha\otimes\beta$ is not necessarily UDA. Admitting that one of $\alpha$ or $\beta$ is pure, $\alpha\otimes\beta$ is shown to be UDA by three composite types of tensor product add2023shen. Generally, it is worth considering the additivity of two mixed UDA states by focusing on the following questions: can the UDA additivity hold without additional conditions? If the answer is negative, which kind of additional conditions of $\alpha$ and $\beta$ make the UDA additivity hold? The purity assumption in add2023shen avoids analyzing the essence of UDA states, whereas this is unavoidable when both states are mixed. By fully characterizing mixed UDA states, the additivity of these UDA states can be further studied. This is the second motivation of this work.
In this work, we study the complete characterization of multipartite mixed UDA states, and show the necessary and sufficient condition under which the additivity of UDA states holds. As the beginning, we introduce some essential properties for arbitrary UDA states in Lemmas 4-7. In detail, we show that the global state is not UDA if its marginals or faces of the range are not UDA, and that two states with the same range, or up to permutations, are either UDA or not at the same time. We also show the range inclusion between a specific state and its compatible states. By these facts, we show the complete characterization of bipartite 1-UDA states in Proposition 8. The necessary and sufficient conditions by which some three- to $n$-partite states are UDA are given in Theorems 14, 20, 22 and Proposition 23. We clarify that the study of $k$-UDA, particularly for small $k$, is motivated by practical considerations. Since the measurement cost, the error of gates and the impact of noise grow with $k$, minimizing resource consumption in local measurements makes $2$-UDA a favorable scenario. Inspired by the proof of preceding facts, a systematic method is derived for the determination of $k$-UDA states in FIG. 1. The UDA additivity of bipartite, three-qubit 2-UDA and some $n$-partite states is further analyzed. For the bipartite UDA states $\alpha$ and $\beta$, $\alpha\otimes\beta$ is UDA if and only if one of $\alpha$ and $\beta$ is a pure product state, shown in Theorem 26. The additivity of three-qubit UDA states is given in Theorem 27, by showing the necessary and sufficient condition. The applications of mixed UDA states and their characterization are shown from the perspectives of tomography, entanglement certification and other tasks.
The rest of this paper is organized as follows. In Sec. II, we clarify some notations and introduce the definitions and known facts. In Sec. III, we characterize multipartite mixed UDA states and establish a systematic method to determine some non-$k$-UDA states. In Sec. IV, we analyze the additivity of mixed UDA states and present the necessary and sufficient conditions. We show the applications of mixed UDA states and their characterization in Sec. V. We conclude in Sec. VI.
II Preliminaries
In this section, we present the notation and the definition of mixed UDA states, along with some known facts used later.
First, we introduce some notations. Suppose $A_{1},A_{2},...,A_{n}$ are $n$ systems associated with the Hilbert spaces ${\cal H}_{A_{1}},{\cal H}_{A_{2}},...,{\cal H}_{A_{n}}$, respectively. We denote a state $\rho$ on ${\cal H}_{A_{1}}\otimes{\cal H}_{A_{2}}\otimes...\otimes{\cal H}_{A_{n}}$ by $\rho_{A_{1}A_{2}...A_{n}}$. We denote the range and kernel of the state $\rho$ by ${\cal R}(\rho)$ and $\ker(\rho)$, respectively. They are the range and kernel of the Hermitian density operator of $\rho$. We have ${\cal R}(\rho)=\ker(\rho^{\dagger})=\ker(\rho)=\mathop{\rm supp}\rho$. Let $\cal S$} be a subset of $\{A_{1},A_{2},...,A_{n}\}$ with $|{\cal S}|=k$ and ${\cal S}^{c}=\{A_{1},A_{2},...,A_{n}\}\backslash{\cal S}$. A $k$-partite marginal of $\rho$ is given by $\rho_{{\cal S}}=\mathop{\rm Tr}_{{\cal S}^{c}}\rho$. A state $\sigma$ is said to be $k$-compatible with $\rho$ if $\sigma$ has the same $k$-partite marginals as $\rho$, i.e. $\sigma_{{\cal S}}=\rho_{{\cal S}}$ for any $\cal S$}.
Next we introduce the definition of a state uniquely determined by its $k$-partite marginals. For a pure state $|\psi\rangle$, if there is no other pure state $k$-compatible with $|\psi\rangle$, then $|\psi\rangle$ is said to be $k$-partite uniquely determined among pure states ($k$-UDP); if there is no other (pure or mixed) state $k$-compatible with $|\psi\rangle$, then $|\psi\rangle$ is said to be $k$-partite uniquely determined among all ($k$-UDA) states. This definition can be generalized to the case of mixed states by assuming that the initial state $\rho$ is mixed.
Definition 1
Let $\rho$ be an arbitrary state (pure or mixed). If there is no other state having all the same $k$-partite marginals as $\rho$, then $\rho$ is called $k$-uniquely determined among all ($k$-UDA) states.
The existence of mixed $k$-UDA states is established in previous work. It is known that some mixed states with low rank can be UDA. For example, Theorem 4 in uinque2013chen shows that almost every tripartite density operator $\rho$ acting on the Hilbert space ${\cal H}_{d_{1}}\otimes{\cal H}_{d_{2}}\otimes{\cal H}_{d_{3}}$ with rank no more than $\lfloor\frac{d_{1}}{d_{3}}\rfloor$ can be uniquely determined among all states by its 2-marginals of partial pairs $\{1,2\}$ and $\{1,3\}$. It is also shown in uinque2013chen that giving a complete characterization of these undetermined tripartite states is not easy, as the methods proposed before do not directly provide it. This inspires us to give a complete characterization of UDA states before analyzing their additivity. That is, we establish a criterion to determine whether a mixed state is UDA directly, which is exactly what we will show in Sec. III.
A state with fewer parties may transform from a $k$-UDA state into a non-$k$-UDA one, with the increase of the number of parties with a fixed $k\in\mathbb{N}$. A construction of $k$-UDA states in a system with more parties is considered in add2023shen. In detail, three composite types of tensor product for two $k$-UDA states preserve $k$-UDA if one of the two $k$-UDA states is pure. This relies on the following fact, which we shall use later.
Lemma 2
If the reduction of the system $(A_{1},\ldots,A_{m})$ from the global state $\rho_{A_{1}\cdots A_{m}E}$ is a pure state, then the global state is in the form
$\rho_{A_{1}\cdots A_{m}E}=|\psi\rangle\!\langle\psi|_{A_{1}\cdots A_{m}}\otimes\sigma_{E}.$ (1)
Obviously, Lemma 2 does not generally hold if the $m$-partite marginal $\rho_{A_{1}\cdots A_{m}}$ is a mixed state. For example, the state $\rho_{ABC}=p|000\rangle\!\langle 000|+(1-p)|W\rangle\!\langle W|$ with $|W\rangle=\frac{1}{\sqrt{3}}(|000\rangle+|001\rangle+|100\rangle)$ satisfies that $\rho_{AB}$ and $\rho_{C}$ are both mixed, and $\rho_{ABC}\neq\rho_{AB}\otimes\rho_{C}$. Hence, the construction in add2023shen fails if both initial states are mixed. It is interesting to establish additional conditions on the initial states, by which the additivity of mixed UDA states still holds.
We introduce some basic properties for a $k$-UDA state add2023shen as follows. Let $\rho$ be a $n$-partite $k$-UDA state. Then (i) any state that is LU equivalent to $\rho$ is also a $k$-UDA state; (ii) $\rho$ is also $(k+1)$-UDA state for $k+1\leq n$, and the converse does not generally hold. These facts allow us to simplify the form of density operators in the analysis of their UDA properties, which shall be widely used in Sec. III. In addition to the equation of $k$-marginals, the positive semi-definiteness is a natural restriction of a compatible state. To prove that some states are not $k$-UDA, we construct another state compatible with them, and the semi-definiteness of these states is guaranteed by the following fact.
Lemma 3
If $X$ and $Y$ are two positive semidefinite matrices, and ${\cal R}(X)\subseteq{\cal R}(Y)$, then there is a small enough positive number $\varepsilon>0$ such that $Y-\varepsilon X\geq 0$.
Proof.
From $X\geq 0$ and $Y\geq 0$, there is an invertible matrix $P$ such that $P^{\dagger}XP=\begin{bmatrix}D_{1}&0\\ 0&0\end{bmatrix}$ and $P^{\dagger}YP=\begin{bmatrix}D_{2}&0\\ 0&0\end{bmatrix}$, where $D_{1}=\mathop{\rm diag}\{s_{1},s_{2},...,s_{r_{1}}\}$ and $D_{2}=\mathop{\rm diag}\{t_{1},t_{2},...,t_{r_{2}}\}$ are diagonal matrices with positive elements. By ${\cal R}(X)\subseteq{\cal R}(Y)$, we have $\mathop{\rm rank}(D_{1})\leq\mathop{\rm rank}(D_{2})$. Let $\varepsilon=\mathop{\rm min}_{k=1}^{r_{1}}\frac{t_{k}}{s_{k}}$. Then $P^{\dagger}(Y-\varepsilon X)P\geq 0$, i.e. $Y-\varepsilon X\geq 0$.   $\sqcap\sqcup$
Although we may establish different states by Lemma 3, we shall emphasize that the positive semi-definiteness is a power restriction that helps to eliminate the uncertainty of the state and conclude the UDA property, as we will see in Sec. III.3.
III The characterization of multipartite mixed UDA states
In this section, we show the necessary and sufficient conditions with which the bipartite to $n$-partite mixed state is UDA, and establish a systematic method to determine some non-$k$-UDA states. We present the essential properties of UDA states in Sec. III.1. By these properties, we show the necessary and sufficient condition by which a bipartite mixed state is 1-UDA in Sec. III.2. We further characterize the tripartite and four-qubit mixed 2-UDA states in Sec. III.3 and III.4, respectively. We analyze the $n$-qubit mixed $k$-UDA states for $n\geq 5$, and establish a systematic method to determine some non-$k$-UDA states in Sec. III.5. Our complete characterization of tripartite to $n$-partite states mostly focus on 2-UDA states, as such states requires less local information than $k$-UDA states for $k\geq 3$. Hence, 2-UDA states are of practical interest.
III.1 Preliminary lemmas
We show some essential properties satisfied by arbitrary $k$-UDA states from the perspective of their ”components” and range. We also show that the UDA property is invariant to the permutation operations. First, Lemma 4 and 5 show that if the marginals and the element of the range of a state is not $k$-UDA, then the whole state is also not $k$-UDA, respectively. Then we obtain that two states that have the same range are the same time $k$-UDA or not. Lemma 6 shows that the states compatible with each other satisfy range inclusion. Besides, Lemma 7 shows that the UDA property is invariant up to system permutation. The proof of Lemmas 4-7 is given in Appendix A.
Lemma 4
Suppose one of the states $\alpha$ and $\beta$ is not $k$-UDA. Then $\alpha\otimes\beta$ is also not $k$-UDA.
We will use Lemma 4 to conclude that some states with non-$k$-UDA marginals are not 2-UDA, which simplifies the analysis of the additivity of UDA states.
For any subspace $E\subseteq\mathbb{C}^{m}\otimes\mathbb{C}^{n}$, the face of $E$ is defined as the set of all states whose range is contained in $E$. The following fact shows that if the face of ${\cal R}(\rho)$ has a non-$k$-UDA state, then $\rho$ is not $k$-UDA. It is also shown that two states with the same range are $k$-UDA simultaneously, and that two global states are $k$-compatible, implying that their marginals are also $k$-compatible.
Lemma 5
Let $\rho$ and $\sigma$ be two $n$-partite states. Then
(i) If $\rho$ is not $k$-UDA, the convex combination of $\rho$ and $\sigma$, i.e. $\lambda\rho+(1-\lambda)\sigma$ is also not $k$-UDA, for $\lambda\in(0,1]$. Hence, if the face of ${\cal R}(\rho)$ has a non-$k$-UDA state, then $\rho$ is not $k$-UDA.
(ii) If $\rho$ and $\sigma$ have the same range, then they are at the same time k-UDA or not.
(iii) Suppose $\rho$ and $\sigma$ are on ${\cal H}_{A_{1}A_{2}...A_{n}}$ are $k$-compatible with $n>k$. Let ${\cal S}\subseteq(A_{1},...,A_{n})$ such that $|{\cal S}|\geq k$. Then $\rho_{\cal S}$ and $\sigma_{\cal S}$ are also $k$-compatible.
Lemma 5 (i) inspires us to construct the non-$k$-UDA states systematically, say if a state $\alpha$ is not $k$-UDA, any mixed state $\rho$ whose convex combination includes $\alpha$ is also not 2-UDA. It also shown that the states in the face of a $k$-UDA state are also $k$-UDA, which also derives the establishment of $k$-UDA states. Besides, Lemma 5 (ii) allows us to simplify the expression of states by keeping their range invariant. If $\alpha$ is not $k$-UDA and ${\cal R}(\alpha)\subseteq{\cal R}(\beta)$, then $\beta-\epsilon\alpha$ is also not $k$-UDA, for $\epsilon>0$ small enough.
It is vital to establish the relation of a state with all states compatible with it. Here, we consider the relation by the range of the states. We show that two specific product states that are compatible satisfy the range inclusion condition.
Lemma 6
Let $\rho=\rho_{A_{1}A_{2}...A_{n}}\otimes\rho_{A_{n+1}A_{n+2}...A_{m}}$ be a $m$-partite state for $m\geq 2n$. Suppose $\rho_{A_{n+1}A_{n+2}...A_{m}}$ is $n$-UDA. Then the $m$-partite state $\sigma$ $n$-compatible with $\rho$ satisfies ${\cal R}(\sigma)\subseteq{\cal R}(\rho)$.
The operator $\chi=\rho-\sigma$ for any $\sigma$ satisfies that any $n$-marginals are zero if $\sigma$ is compatible with $\rho$. One can verify that $\rho$ is $n$-UDA if and only if $\chi=0$ for any $\sigma$. Lemma 6 implies that such operator $\chi$ satisfies that ${\cal R}(\chi)\subseteq{\cal R}(\rho)$, by which we show that $\chi=0$ for some $\rho$.
Generally, the UDA property can only be invariant up to LU equivalence due to the partial trace invariance of local unitaries. A special nonlocal operation, such as a permutation, can also preserve the UDA property of states. This gives the following fact.
Lemma 7
Two states $\alpha\otimes\beta$ and $\beta\otimes\alpha$ are the same time $k$-UDA or not.
The preceding facts contribute to simplifying the analysis of UDA properties by applying permutation operations. In this paper, we shall interchange the order of any two subsystems of a product state without explicit clarification.
III.2 On bipartite mixed UDA states
We show the necessary and sufficient condition by which a bipartite mixed state is 1-UDA. The additivity of 1-UDA states $\alpha$ and $\beta$ is derived by this characterization.
Proposition 8
A bipartite mixed state $\rho$ on systems $A_{1}$ and $A_{2}$ is 1-UDA if and only if one of $\rho_{A_{1}}$ and $\rho_{A_{2}}$ is pure. Hence if $\rho_{A_{2}A_{3}}$ is 1-UDA then $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}$ is 2-UDA.
Proof.
The if part is trivial by Lemma 2. Next, we show the ”only if” part. Let $\gamma_{A_{1}A_{2}}$ be a bipartite mixed state. The UDA property is invariant up to local unitaries. It suffices to discuss the UDA property of $\rho_{A_{1}A_{2}}=(U\otimes V)\gamma_{A_{1}A_{2}}(U^{\dagger}\otimes V^{\dagger})$. Note that
$\rho_{A_{1}}=U\gamma_{A_{1}}U^{\dagger},\quad\rho_{A_{2}}=V\gamma_{A_{2}}V^{\dagger}.$ (2)
We can choose appropriate local unitaries $U$ and $V$, such that $\rho_{A_{1}}$ and $\rho_{A_{2}}$ are both diagonal. Let $\mathop{\rm Dim}({\cal H}_{A_{k}})=d_{k}$ for $k=1,2$. Then $\rho_{A_{1}A_{2}}$ can be given by $\rho_{A_{1}A_{2}}=\sum_{i,j=1}^{d_{1}}|i\rangle\!\langle j|\otimes M_{ij}$, where $M_{ij}\in\mathbb{C}^{d_{2}\times d_{2}}$, and
$\rho_{A_{1}}=\sum_{i,j=1}^{d_{1}}|i\rangle\!\langle j|\otimes\mathop{\rm Tr}M_{ij},\quad\rho_{A_{2}}=\sum_{j=1}^{d_{1}}M_{jj},$ (3)
are both diagonal states. We further assume that $\rho_{A_{1}A_{2}}=(r_{jk})$ for $j,k=1,2,...,d_{1}d_{2}$. If there is an element $r_{jk}\neq 0$ for $j\neq k$, then $\rho_{A_{1}A_{2}}$ is not 1-UDA. In fact, we can choose the state $\sigma^{\prime}_{A_{1}A_{2}}$ whose diagonal elements are the same as those of $\rho$, and off-diagonal elements are half of those of $\rho$, that is,
$\sigma^{\prime}=\begin{bmatrix}r_{11}&\frac{r_{12}}{2}&\cdots&\frac{r_{1,d_{1}d_{2}}}{2}\\ \frac{r_{12}^{*}}{2}&r_{22}&\cdots&\frac{r_{2,d_{1}d_{2}}}{2}\\ \vdots&\vdots&&\vdots\\ \frac{r_{1,d_{1}d_{2}}^{*}}{2}&\frac{r_{2,d_{1}d_{2}}^{*}}{2}&\cdots&r_{d_{1}d_{2},d_{1}d_{2}}\end{bmatrix}$. (4)
We have $\sigma^{\prime}\neq\rho$ and $\sigma^{\prime}\geq 0$ by $\sigma^{\prime}-\frac{1}{2}\rho\geq 0$. From (3), one can vefify that $\sigma^{\prime}_{A_{k}}=\rho_{A_{k}}$, i.e., $\sigma^{\prime}$ is compatible with $\rho$ by two 1-marginals. Hence, a 1-UDA bipartite state should be diagonal.
Next we show that $\rho_{A_{1}A_{2}}$ is 1-UDA implies that $\rho_{A_{1}}$ or $\rho_{A_{2}}$ is a pure state. Otherwise it holds that $\mathop{\rm rank}(\rho_{A_{k}})\geq 2$. Since $\rho_{A_{k}}$’s are diagonal, there is another state $\sigma_{1}=\eta\otimes\xi$ that is compatible with $\rho$, where $\eta=\rho_{A_{1}}:=\mathop{\rm diag}\{a_{1},a_{2},...,a_{d_{1}}\}$ and $\xi=\rho_{A_{2}}:=\mathop{\rm diag}\{b_{1},b_{2},...,b_{d_{2}}\}$. Up to local unitaries, we assume that $a_{1}\geq a_{2}>0$ and $b_{1}\geq b_{2}>0$. There is a $0<\varepsilon<\mathop{\rm min}\{a_{1},a_{2},b_{1},b_{2}\}$. The state $\sigma_{1}$ can be decomposed as
$\sigma_{1}=(\alpha_{1}+\alpha_{2})\otimes(\beta_{1}+\beta_{2}),$ (5)
where $\alpha_{1}=\beta_{1}=\mathop{\rm diag}\{\varepsilon,\varepsilon,0,...,0\}$, $\alpha_{2}=\mathop{\rm diag}\{a_{1}-\varepsilon,a_{2}-\varepsilon,a_{3},...,a_{d_{1}}\}$ and $\alpha_{3}=\mathop{\rm diag}\{b_{1}-\varepsilon,b_{2}-\varepsilon,b_{3},...,b_{d_{1}}\}$. Note that the state $\alpha_{1}\otimes\beta_{1}$ is compatible with a bipartite entangled state $\tau=2\varepsilon^{2}(|00\rangle+|11\rangle)(\langle 00|+\langle 11|)$. Hence $\sigma_{1}$ is compatible with a state
$\sigma_{2}=\tau+\alpha_{1}\otimes\beta_{2}+\alpha_{2}\otimes(\beta_{1}+\beta_{2}),$ (6)
and $\sigma_{1}\neq\sigma_{2}$. That is, $\rho$ is compatible with two different states $\sigma_{1}$ and $\sigma_{2}$. This shows that $\rho$ is not 1-UDA. This completes the proof.   $\sqcap\sqcup$
It has been shown that if one of $\rho_{A_{1}}$ and $\rho_{A_{2}}$ is pure, then $\rho$ is 1-UDA. Proposition 8 further shows that this is the only possible condition that $\rho$ is 1-UDA. Therefore, the UDA property requires that the rank of the density operator is relatively low.
III.3 On tripartite mixed UDA states
We characterize a three-qubit product-mixed 2-UDA state by deriving the necessary and sufficient condition it satisfies. By this fact, we can determine whether such a state is 2-UDA directly. It is helpful for the analysis of the additivity of mixed UDA states.
As the beginning, we consider a three-qubit state $\rho$ in the form of $\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}$. Obviously, if one of the six 1- and 2-marginals of $\rho$ is pure, then $\rho$ is 2-UDA. In fact, we assume that $\rho_{A_{1}}$ or $\rho_{A_{2}A_{3}}$ is pure. Then $\rho=|\eta\rangle\!\langle\eta|_{A_{1}}\otimes\rho_{A_{2}A_{3}}$ (resp. $\rho=\rho_{A_{1}}\otimes|\psi\rangle\!\langle\psi|_{A_{2}A_{3}}$). By Lemma 2, we obtain that a state $\sigma$ that is 2-compatible with $\rho$ satisfies that $\sigma=|\eta\rangle\!\langle\eta|_{A_{1}}\otimes\sigma_{A_{2}A_{3}}$ (resp. $\sigma=\sigma_{A_{1}}\otimes|\psi\rangle\!\langle\psi|_{A_{2}A_{3}}$). By $\rho_{A_{2}A_{3}}=\sigma_{A_{2}A_{3}}$ (resp. $\rho_{A_{1}A_{2}}=\sigma_{A_{1}A_{2}}$), we have $\rho=\sigma$. We show that this condition is necessary and sufficient for a state $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}$ with diagonal $\rho_{A_{2}A_{3}}$.
Proposition 9
Suppose $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}$ is a tripartite state and $\rho_{A_{2}A_{3}}$ is diagonal. Then $\rho$ is 2-UDA if and only if one of the six 1- and 2-marginals of $\rho$ is pure.
The proof of Proposition 9 is shown in Appendix B. This implies that a tripartite state $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}}\otimes\rho_{A_{3}}$ is 2-UDA if and only if one of $\rho_{A_{1}}$, $\rho_{A_{2}}$, $\rho_{A_{3}}$ is a pure state by applying appropriate local unitaries.
For a state $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}$ with arbitrary two-qubit state $\rho_{A_{2}A_{3}}$, the result in Proposition 9 does not generally hold. Here is an example.
Example 10
We consider the state $\xi^{(x)}_{A_{1}A_{2}A_{3}}=(\frac{1}{2}I_{2})_{A_{1}}\otimes\xi_{A_{2}A_{3}}^{(x)}$ with $\xi^{(x)}_{A_{2}A_{3}}=\begin{bmatrix}\frac{1}{3}&0&0&0\\ 0&\frac{1}{3}&x&0\\ 0&x^{*}&\frac{1}{3}&0\\ 0&0&0&0\end{bmatrix}$. By choosing $x=1/3$, one can verify that $\xi^{(1/3)}$ is 2-UDA. Besides, we have $\mathop{\rm rank}(\xi^{(1/3)}_{A_{1}A_{j}})=4$ for $j=2,3$ and $\mathop{\rm rank}(\xi^{(1/3)}_{A_{2}A_{3}})=\mathop{\rm rank}(\xi^{(1/3)}_{A_{k}})=2$ for $k=1,2,3$, i.e. neither of the 1- and 2-marginals of $\xi^{(1/3)}$ is pure. Further we have $\xi$ is 2-UDA iff $|x|=\frac{1}{3}$.
Inspired by this example, we consider the necessary and sufficient condition with which the state $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}}$ is 2-UDA, according to the rank of $\rho_{A_{2}A_{3}}$. By preceding analysis, it suffices to consider the case that $\mathop{\rm rank}(\rho_{A_{2}A_{3}})>1$, since a pure $\rho_{A_{2}A_{3}}$ implies that $\rho$ is 2-UDA.
The UDA property is invariant up to local unitaries. It allows us to simplify a bipartite state as follows.
Lemma 11
Let $\eta$ be a bipartite state supported on $\mathbb{C}^{m}\otimes\mathbb{C}^{n}$ and $\mathop{\rm rank}(\eta)=r$ with $1\leq r\leq m+n-2$. Up to local unitaries, the last diagonal element of $\eta$ is zero, i.e. $\eta=M\oplus 0$ with $M\geq 0$ and $\mathop{\rm Tr}(M)=1$.
Proof.
The dimension $d$ of the kernel of $\eta$ satisfies that $d=mn-r\geq(m-1)(n-1)+1$. By the theory of Segre variety, there is a product state in the kernel of $\eta$. Up to local unitaries, we assume that this product is $|m-1,n-1\rangle$. Then the range of $\eta$ does not contain $|m-1,n-1\rangle$, and thus $\eta=M\oplus 0$.   $\sqcap\sqcup$
By choosing $m,n,r=2$ in Lemma 11, we show the necessary and sufficient condition that a set of product mixed states ${\cal P}=\{\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}:\mathop{\rm rank}(\rho_{A_{2}A_{3}})=2\text{ or }\mathop{\rm rank}(\rho_{A_{2}A_{3}})=3\text{ with a zero diagonal element}\}$ is 2-UDA.
Proposition 12
Suppose a three-qubit state $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}$ with
$\rho_{A_{2}A_{3}}=D\oplus 0\text{ for }D=\begin{bmatrix}q_{1}&y&z\\ y^{*}&q_{2}&x\\ z^{*}&x^{*}&q_{3}\end{bmatrix}$. (7)
Then (i) if $\mathop{\rm rank}(\rho_{A_{1}})=1$ or $q_{2}=0$ or $q_{3}=0$, then $\rho$ is 2-UDA.
(ii) if $\mathop{\rm rank}(\rho_{A_{1}})=2$ and $q_{k}>0$, then $\rho$ is 2-UDA iff $|D|=0$.
(iii) if $\mathop{\rm rank}(\rho_{A_{1}})=2$ and $q_{1}=0$, then $\rho$ is 2-UDA iff $|x|=\sqrt{q_{2}q_{3}}$.
Proposition 12 shows that the state $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}$ with a rank-three $\rho_{A_{2}A_{3}}$ which has a zero diagonal element is not 2-UDA. The following fact is given for a high-dimensional state. By this fact, one can obtain that $\rho$ is not 2-UDA if $\mathop{\rm rank}(\rho_{A_{1}})=2$ and $\mathop{\rm rank}(\rho_{A_{2}A_{3}})=3$ or 4, by choosing $d_{1}=d_{2}=d_{2}=2$.
Proposition 13
Let $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}$ be a tripartite state supported on $\mathbb{C}^{d_{1}}\otimes\mathbb{C}^{d_{2}}\otimes\mathbb{C}^{d_{3}}$. Suppose $\mathop{\rm rank}(\rho_{A_{1}})=d_{1}$. Then
(i) If $\mathop{\rm rank}(\rho_{A_{2}A_{3}})=d_{2}d_{3}-1$ and $\rho_{A_{2}A_{3}}$ is not LU equivalent to the state with a zero diagonal element, then $\rho$ is not 2-UDA.
(ii) If $\mathop{\rm rank}(\rho_{A_{2}A_{3}})=d_{2}d_{3}$, then $\rho$ is not 2-UDA.
The proof of Proposition 13 is given in Appendix B.
By summarizing the facts in Propositions 12 and 13, we obtain the complete characterization of a 2-UDA product state $\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}$ as follows.
Theorem 14
Let $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}$ be a three-qubit state. Then $\rho$ is 2-UDA if and only if $\rho$ satisfies one of the following conditions,
(i) at least one of the 1-marginals of $\rho$ is pure;
(ii) $\rho_{A_{2}A_{3}}$ is pure;
(iii) $\mathop{\rm rank}(\rho_{A_{1}})=\mathop{\rm rank}(\rho_{A_{2}A_{3}})=2$ and $\rho_{A_{2}A_{3}}$ is LU equivalent with $D\oplus 0$ in (7) with $q_{1}q_{2}q_{3}>0$.
Proof.
First, we claim that if $\mathop{\rm rank}(\rho_{A_{1}})=1$ or $\mathop{\rm rank}(\rho_{A_{2}A_{3}})=1$, then $\rho$ is 2-UDA. The proof is trivial by Lemma 2.
Second, if $\mathop{\rm rank}(\rho_{A_{1}})=2$ and $\mathop{\rm rank}(\rho_{A_{2}A_{3}})=2$, then $\rho$ is LU-equivalent to $\rho_{A_{1}}\otimes(D\oplus 0)_{A_{2}A_{3}}$ in (7). If $q_{1}q_{2}q_{3}>0$ in (7) or $q_{2}=0$ or $q_{3}=0$ in $D$, then $\rho$ is 2-UDA; If $q_{1}=0$, then $\rho$ is not 2-UDA. The first claim is derived by Proposition 12. Next, If $q_{2}=0$ or $q_{3}=0$ in $D$, then $\rho$ is 2-UDA by Proposition 12 (i). if $q_{k}>0$, then $\mathop{\rm rank}(\rho_{A_{2}A_{3}})=2$ implies that $|D|=0$, and $\rho$ is 2-UDA by Proposition 12 (ii). If $q_{1}=0$ and $\mathop{\rm rank}(\rho_{A_{2}A_{3}})=2$, then $|x|<\sqrt{q_{2}q_{3}}$, and $\rho$ is not 2-UDA by Proposition 12 (iii).
Finally, if $\mathop{\rm rank}(\rho_{A_{1}})=2$ and $\mathop{\rm rank}(\rho_{A_{2}A_{3}})=3\text{ or }4$ , then $\rho$ is not 2-UDA. In fact, if $\rho_{A_{2}A_{3}}$ is LU equivalent to a state that has a zero diagonal element, all the remaining diagonal elements are strictly larger than zero, and $\rho$ is not 2-UDA by Proposition 12 (ii). Otherwise, $\rho$ is not 2-UDA by Proposition 13 (i). The result for full-rank $\rho_{A_{2}A_{3}}$ is directly obtained by Proposition 13 (ii).   $\sqcap\sqcup$
Based on Theorem 14, the whole three-qubit mixed states that satisfy one of the three conditions can be determined by their 2-marginals. This implies that it is practically redundant to perform tomography on the system $A_{1}A_{2}A_{3}$.
III.4 On four-qubit mixed UDA states
By virtue of the essential characterization of the three-qubit 2-UDA state, we present the necessary and sufficient condition by which a mixed state $\rho=\rho_{A_{1}A_{2}}\otimes\rho_{A_{3}A_{4}}$ is 2-UDA. It is shown that it is possible for the state to be 2-UDA only when the rank of $\rho$ is less than six. We further deal with the ”low rank” case and show the necessary and sufficient condition in Theorem 20.
As the first step, we consider the diagonal 2-UDA state and give its characterization.
Proposition 15
Let $\rho=\rho_{A_{1}A_{2}}\otimes\rho_{A_{3}A_{4}}$ be a four-qubit diagonal state. Then $\rho$ is 2-UDA if and only if one of the following conditions holds, (i) $\rho_{A_{1}A_{2}}$ or $\rho_{A_{3}A_{4}}$ is pure; (ii) at least two of the 1-marginals of $\rho$ are pure.
Proof.
The ”if part” is trivial. We show the ”only if” part. Suppose $\rho_{A_{1}A_{2}}$, $\rho_{A_{3}A_{4}}$ are both mixed, and there is at most one 1-marginal that is pure. Without loss of generality, we assume that the 1-marginal $\rho_{A_{1}}$ is pure merely. Then $\rho$ is 2-UDA iff $\rho_{A_{2}}\otimes\rho_{A_{3}A_{4}}$ is 2-UDA. From Theorem 14, $\rho$ is not 2-UDA. Otherwise, none of the 1- and 2-marginals of $\rho$ is pure. One can establish another state $\sigma$ which has a nonzero element, and is compatible with $\rho$. This completes the proof.   $\sqcap\sqcup$
As we have done before, we consider the tensor product of rank-two two-qubit states. In this case, the state may or may not be 2-UDA, depending on the form of the state. So it deserves detailed analysis. From Lemma 11, up to LU equivalence, we can assume that
$\rho=\rho_{A_{1}A_{2}}\otimes\rho_{A_{3}A_{4}}:=\begin{bmatrix}p_{1}&b&c&0\\ b^{*}&p_{2}&a&0\\ c^{*}&a^{*}&p_{3}&0\\ 0&0&0&0\end{bmatrix}\otimes\begin{bmatrix}q_{1}&y&z&0\\ y^{*}&q_{2}&x&0\\ z^{*}&x^{*}&q_{3}&0\\ 0&0&0&0\end{bmatrix}$. (8)
First, we assume that $\rho_{A_{1}A_{2}}$ is a product state, and analyze its 2-UDA property. From Lemma 4, we have the following fact. If the four-qubit state $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}}\otimes\rho_{A_{3}A_{4}}$ is 2-UDA, then $\rho_{A_{2}}\otimes\rho_{A_{3}A_{4}}$ and $\rho_{A_{1}}\otimes\rho_{A_{3}A_{4}}$ is 2-UDA. It is analyzed in Theorem 14. If one of $\rho_{A_{1}}$ and $\rho_{A_{2}}$ is pure, then $\rho$ is 2-UDA iff $\rho_{A_{k}}\otimes\rho_{A_{3}A_{4}}$ is 2-UDA for $k=1,2$. We assume that $\rho_{A_{k}}$ is mixed in the subsequent analysis. From Theorem 14, $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}}\otimes\rho_{A_{3}A_{4}}$ is 2-UDA implies that one of the three conditions holds, (a) $\rho_{A_{3}}$ or $\rho_{A_{4}}$ is pure, i.e. $\rho=\otimes_{k=1}^{4}\rho_{A_{k}}$. Further Proposition 15 implies that $\rho_{A_{3}}$ and $\rho_{A_{4}}$ are both pure since $\rho_{A_{1}}$ and $\rho_{A_{2}}$ are mixed. (b) $\rho_{A_{3}A_{4}}$ is pure. (c) $\mathop{\rm rank}(\rho_{A_{3}A_{4}})=2$ and $\rho_{A_{3}A_{4}}$ is LU equivalent to $D\oplus 0$ with the positive diagonal elements of $D$.
Note that (a) and (b) imply that $\rho$ is 2-UDA. We show that (c) is not a sufficient condition of 2-UDA $\rho$.
Lemma 16
Suppose $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}}\otimes\rho_{A_{3}A_{4}}$ is a four-qubit state. Here $\mathop{\rm rank}(\rho_{A_{1}})=\mathop{\rm rank}(\rho_{A_{2}})=\mathop{\rm rank}(\rho_{A_{3}A_{4}})=2$ and $\rho_{A_{3}A_{4}}$ is LU equivalent to $D\oplus 0$ in (7) with the positive diagonal elements of $D$. Then $\rho$ is not 2-UDA.
Proof.
Since $\rho_{A_{1}}\otimes\rho_{A_{2}}$ is full-rank, there is $D=\mathop{\rm diag}\{0,\epsilon_{1},\epsilon_{2},0\}$ with $\epsilon_{1},\epsilon_{2}$ small enough, such that $\xi:=\rho_{A_{1}}\otimes\rho_{A_{2}}-D\geq 0$ by Lemma 3. Then $\rho=\xi\otimes\rho_{A_{3}A_{4}}+D\otimes\rho_{A_{3}A_{4}}$. We show that
$\alpha:=D\otimes\rho_{A_{3}A_{4}}$ (9)
is not 2-UDA. From Lemma 5 (ii), it suffices to consider the range of $\rho_{A_{3}A_{4}}$, up to LU equivalence, is given as follows, for $zuv\neq 0$ and $nh\neq 0$,
${\cal R}(\rho_{A_{3}A_{4}})=\text{span}\{z|00\rangle,\;u|01\rangle+v|10\rangle\},$ (10)
${\cal R}(\rho_{A_{3}A_{4}})=\text{span}\{m|00\rangle+n|01\rangle,\;g|00\rangle+h|10\rangle\}.$ (11)
Suppose $\sigma$ is a four-qubit state compatible with $\alpha$. Then
$\sigma=0_{4\times 4}\oplus H\oplus 0_{4\times 4},$ (12)
where $H=\begin{bmatrix}\epsilon_{1}\rho_{A_{3}A_{4}}&B\\ B^{\dagger}&\epsilon_{2}\rho_{A_{3}A_{4}}\end{bmatrix}$ for $B=B_{1}\oplus 0$ and $\mathop{\rm Tr}{B}=0$. By some observations, we can always choose appropriate $B_{1}$ such that $\lVert B\rVert$ is small enough and
${\cal R}(B^{\dagger})\subseteq{\cal R}(\rho_{A_{3}A_{4}})$ (13)
${\cal R}(BB^{\dagger})={\cal R}(B)\subseteq{\cal R}(\rho_{A_{3}A_{4}}),$ (14)
where (14) is implied by (13). In detail, we consider the case for the range of $\rho_{A_{3}A_{4}}$ is given in (10). We choose $B_{1}=\begin{bmatrix}-(y_{1}^{2}+y_{2}^{2})&0&0\\ 0&y_{1}^{2}&y_{1}y_{2}\\ 0&y_{1}y_{2}&y_{2}^{2}\end{bmatrix}$ for $y_{i}>0$ small enough and $\frac{y_{1}}{y_{2}}=\frac{u}{v}$. Second, the range of $\rho_{A_{3}A_{4}}$ is given as in (11). For $mg\neq 0$, we choose $B_{1}=\frac{1}{\mu}\begin{bmatrix}x_{1}^{2}x_{4}^{2}-x_{2}^{2}x_{3}^{2}&x_{1}x_{2}(x_{3}^{2}+x_{4}^{2})&-x_{3}\mu\\ x_{1}x_{2}(x_{3}^{2}+x_{4}^{2})&x_{2}^{2}(x_{3}^{2}+x_{4}^{2})&0\\ -x_{3}\mu&0&-\,x_{4}\mu\end{bmatrix}$ for $\mu=x_{4}(x_{1}^{2}+x_{2}^{2})$, $x_{i}>0$ small enough, $\frac{x_{1}}{x_{2}}=\frac{m}{n}$ and $\frac{x_{3}}{x_{4}}=\frac{g}{h}$.
Since $\mathop{\rm rank}(\rho_{A_{3}A_{4}})=2$, there is a unitary $V$ such that $H$ is congruent to $\begin{bmatrix}\epsilon_{1}\rho_{A_{3}A_{4}}&BV\\ V^{\dagger}B^{\dagger}&I_{2}\oplus 0_{2}\end{bmatrix}$. Using (13), we obtain that ${\cal R}(BV)\subset{\cal R}(I_{2}\oplus 0_{2})$, and thus $H$ is also congruent to
$(\epsilon_{2}\rho_{A_{3}A_{4}}-BB^{\dagger})\oplus I_{2}\oplus 0_{2}.$
Using Lemma 3 and (14), for $x_{i},y_{i}$ small enough, we have $(\epsilon_{2}\rho_{A_{3}A_{4}}-BB^{\dagger})\geq 0$, and hence $H\geq 0$. Then $\sigma\geq 0$. This shows that $\alpha$ is not 2-UDA. Using Lemma 5, $\rho$ is not 2-UDA.   $\sqcap\sqcup$
Using Lemma 16, $\rho$ is 2-UDA iff (a) or (b) holds. In summary, we have the following fact.
Proposition 17
Let $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}}\otimes\rho_{A_{3}A_{4}}$ be a four-qubit state. Then
(i) If at least one of $\rho_{A_{1}}$ or $\rho_{A_{2}}$ is pure, then $\rho$ is 2-UDA iff the remaining three-qubit state $\rho_{A_{k}}\otimes\rho_{A_{3}A_{4}}$ is 2-UDA given in Theorem 14.
(ii) If both $\rho_{A_{1}}$ and $\rho_{A_{2}}$ are mixed, then $\rho$ is 2-UDA iff one of $\rho_{A_{3}A_{4}}$ is pure.
We show the characterization of the 2-UDA state $\rho_{A_{1}A_{2}}\otimes\rho_{A_{3}A_{4}}$ with $\mathop{\rm rank}(\rho_{A_{1}A_{2}})=\mathop{\rm rank}(\rho_{A_{3}A_{4}})=2$. In this case, $\rho$ may or may not be 2-UDA, depending on the specific form of the density operators. We will know that this is the critical threshold determining whether a quantum state can be 2-UDA, since if a state with the lower (resp. higher) rank will be definitely 2-UDA (resp. not 2-UDA).
Proposition 18
Let $\rho=\rho_{A_{1}A_{2}}\otimes\rho_{A_{3}A_{4}}$ be a four-qubit state in (8) with $\mathop{\rm rank}(\rho_{A_{1}A_{2}})=\mathop{\rm rank}(\rho_{A_{3}A_{4}})=2$. Then
(i) If $p_{2}p_{3}=0$, then $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}}\otimes\rho_{A_{3}A_{4}}$ has been investigated in Proposition 17.
(ii) If $p_{2}p_{3}>0$ and $p_{1}=0$ or $q_{2}q_{3}>0$ and $q_{1}=0$, $\rho$ is not 2-UDA.
(iii) If $p_{1}p_{2}p_{3}>0$ and $q_{1}q_{2}q_{3}>0$, then $\rho$ is 2-UDA.
The proof of Proposition 18 is shown in Appendix C.
The UDA-property of a high-rank product state is analyzed by the results in Proposition 18. This shows that a mixed state with a ”relatively high” rank is less likely to be 2-UDA.
Proposition 19
If one of $\mathop{\rm rank}(\rho_{A_{1}A_{2}})$ and $\mathop{\rm rank}(\rho_{A_{3}A_{4}})$ has rank not less than three, and the other is not pure, then $\rho=\rho_{A_{1}A_{2}}\otimes\rho_{A_{3}A_{4}}$ is not 2-UDA.
The proof of Proposition 19 is presented in Appendix C. The UDA property is closely related to the state’s rank. Proposition 19 inspires us to verify the rank of a mixed state before we determine whether it is 2-UDA. For the high-rank states $\rho$, the tomography performed on ${\cal H}_{A_{1}A_{2}A_{3}A_{4}}$ is necessary.
By now, we have completed the analysis of the 2-UDA property of any four-qubit state $\rho=\rho_{A_{1}A_{2}}\otimes\rho_{A_{3}A_{4}}$. We conclude the results of Propositions 17, 18, and 19, and obtain the necessary and sufficient condition by which $\rho$ is 2-UDA.
Theorem 20
Let $\rho=\rho_{A_{1}A_{2}}\otimes\rho_{A_{3}A_{4}}$ be a four-qubit state. Then $\rho$ is 2-UDA if and only if one of the following conditions holds,
(i) $\rho_{A_{1}A_{2}}$ or $\rho_{A_{3}A_{4}}$ is pure;
(ii) at least one of $\{\rho_{A_{2}},\rho_{A_{3}}\}$ and $\{\rho_{A_{4}},\rho_{A_{5}}\}$ are pure, respectively;
(iii) only one of the 1-marginals $\rho_{A_{j}}$ is pure, and $\rho=\rho_{A_{j}}\otimes\rho_{A_{n}}\otimes\rho_{A_{m_{1}}A_{m_{2}}}$ satisfies that $\mathop{\rm rank}(\rho_{A_{m_{1}}A_{m_{2}}})=2$ and $\rho_{A_{m_{1}}A_{m_{2}}}$ is LU equivalent with $D\oplus 0$ in (7) with $q_{1}q_{2}q_{3}>0$.
(iv) $\rho_{A_{1}A_{2}}$ and $\rho_{A_{3}A_{4}}$ are rank-two, and their product is LU equivalent to
$(H\oplus 0)_{A_{2}A_{3}}\otimes(D\oplus 0)_{A_{4}A_{5}}$ (15)
$=$ $\begin{bmatrix}p_{1}&b&c&0\\ b^{*}&p_{2}&a&0\\ c^{*}&a^{*}&p_{3}&0\\ 0&0&0&0\end{bmatrix}\otimes\begin{bmatrix}q_{1}&y&z&0\\ y^{*}&q_{2}&x&0\\ z^{*}&x^{*}&q_{3}&0\\ 0&0&0&0\end{bmatrix}$,
where $p_{1}p_{2}p_{3}>0$ and $q_{1}q_{2}q_{3}>0$.
As we know, it is difficult to determine whether a state is 2-UDA, even if the state is given numerically. Using Theorem 20, one can verify whether a state is 2-UDA directly. It contributes to analyzing the properties of UDA states, including additivity, thereby making quantum tomography more efficient.
III.5 On $n$-qubit mixed UDA states for $n\geq 5$
We show the necessary and sufficient condition by which some $n$-qubit states are $k$-UDA. We introduce a systematic method for the determination of non-$k$-UDA states in FIG. 1, which derives the construction of multipartite states that is not $k$-UDA.
As the beginning, we show the complete characterization of a 2-UDA five-qubit state $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}}$. The following fact excludes a set of states that is 2-UDA by establishing different states that are compatible with them.
Lemma 21
Suppose $\rho_{A_{1}}$ is a mixed one-qubit state, the two-qubit states $\rho_{A_{2}A_{3}}$ and $\rho_{A_{4}A_{5}}$ satisfy the condition in Theorem 20 (iv). Then the five-qubit state $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}}$ is not 2-UDA.
The proof of Lemma 21 is given in Appendix D. Using Theorems 14 and 20, Proposition 23 and Lemma 21, we have the complete characterization of a five-qubit product state.
Theorem 22
The five-qubit state $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}}$ is 2-UDA if and only if one of the following conditions holds,
(i) $\rho_{A_{2}A_{3}}$ or $\rho_{A_{4}A_{5}}$ is pure, i.e. $\rho=\rho_{A_{1}}\otimes|\psi\rangle\!\langle\psi|_{A_{n_{1}}A_{n_{2}}}\otimes\rho_{A_{m_{1}}A_{m_{2}}}$, and the remaining state $\rho_{A_{1}}\otimes\rho_{A_{m_{1}}A_{m_{2}}}$ satisfies the conditions in Theorem 14.
(ii) $\rho_{A_{1}}$ is pure, and at least one of $\{\rho_{A_{2}},\rho_{A_{3}}\}$ and $\{\rho_{A_{4}},\rho_{A_{5}}\}$ are pure, respectively.
(iii) Up to LU equivalence, $\rho=|\psi\rangle\!\langle\psi|_{A_{1}}\otimes|\phi\rangle\!\langle\phi|_{A_{n_{1}}}\otimes\rho_{A_{n_{2}}}\otimes(D\oplus 0)_{A_{m_{1}}A_{m_{2}}}$, where $n_{1},n_{2}\in\{2,3\}$ or $\{4,5\}$, $\rho_{A_{n_{2}}}$ is rank-two, and $D$ is given in (7).
(iv) Up to LU equivalence, $\rho=|\psi\rangle\!\langle\psi|_{A_{1}}\otimes(H\oplus 0)_{A_{2}A_{3}}\otimes(D\oplus 0)_{A_{4}A_{5}}$, where $H$ and $D$ are given in (15).
Proof.
Obviously, $\rho$ is 2-UDA only if $\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}}$ is 2-UDA by Theorem 20. This derives cases (i)-(iv) as follows. (i) $\rho_{A_{2}A_{3}}$ or $\rho_{A_{4}A_{5}}$ is pure. Then $\rho$ is 2-UDA iff the remaining part is 2-UDA by Lemma 2. (ii) At least one of $\{\rho_{A_{2}},\rho_{A_{3}}\}$ and $\{\rho_{A_{4}},\rho_{A_{5}}\}$ are pure, respectively. Then $\rho=\rho_{A_{1}}\otimes|\psi\rangle\!\langle\psi|_{A_{n_{1}}}\otimes\rho_{A_{n_{2}}}\otimes|\phi\rangle\!\langle\phi|_{A_{m_{1}}}\otimes\rho_{A_{m_{2}}}$ with mixed $\rho_{A_{n_{2}}}$ and $\rho_{A_{m_{2}}}$. Using Theorem 20, $\rho$ is 2-UDA iff $\rho_{A_{1}}\otimes\rho_{A_{n_{2}}}\otimes\rho_{A_{m_{2}}}$ is 2-UDA, i.e. $\rho_{A_{1}}$ is pure. (iii) Up to LU equivalence, $\rho_{A_{2}...A_{5}}=|\phi\rangle\!\langle\phi|_{A_{n_{1}}}\otimes\rho_{A_{n_{2}}}\otimes(D\oplus 0)_{A_{m_{1}}A_{m_{2}}}$. $\rho$ is 2-UDA iff $\rho_{A_{1}}$ is pure by Theorem 20. Otherwise, $\rho_{A_{1}}\otimes\rho_{A_{n_{2}}}\otimes(D\oplus 0)_{A_{m_{1}}A_{m_{2}}}$ is not 2-UDA. (iv) Up to LU equivalence, $\rho_{A_{2}...A_{5}}=(D\oplus 0)_{A_{2}A_{3}}\otimes(H\oplus 0)_{A_{4}A_{5}}$. By Lemma 21, $\rho$ is 2-UDA iff $\rho_{A_{1}}$ is pure.   $\sqcap\sqcup$
We characterize $n$-qubit mixed UDA states. First, we consider the fully product state $\otimes_{i=1}^{n}\rho_{A_{i}}$. On the basis of that, we consider two general cases in (ii) and (iii).
Proposition 23
(i) The $n$-partite state $\rho=\otimes_{i=1}^{n}\rho_{A_{i}}$ is $k$-UDA if and only if at least nkn-k of the 1-marginals $\rho_{A_{i}}$ is pure.
(ii) The $n$-qubit state $\rho=\otimes_{i=1}^{n-2}\rho_{A_{i}}\otimes\rho_{A_{n-1}A_{n}}$ is $2$-UDA if and only if one of the following conditions holds, (a) $\rho_{A_{n-1}A_{n}}$ is pure, and $\rho=\otimes_{i=1}^{n-2}\rho_{A_{i}}$ satisfies (i); (b) $\rho_{A_{n-1}}$ or $\rho_{A_{n}}$ is pure, then $\rho=\otimes_{i=1}^{n}\rho_{A_{i}}$, which satisfies (i); (c) $\otimes_{i=1}^{n-2}\rho_{A_{i}}$ is pure, and $\rho_{A_{n-1}A_{n}}$ is arbitrary; (d) at least n3n-3 of $\rho_{A_{i}}$ is pure, $\rho_{A_{n-1}A_{n}}$ is rank-two and LU equivalent with $D\oplus 0$ in (7) with $q_{1}q_{2}q_{3}>0$.
(iii) The $n$-qubit state $\rho=\otimes_{i=1}^{n-4}\rho_{A_{i}}\otimes\rho_{A_{n-3}A_{n-2}}\otimes\rho_{A_{n-1}A_{n}}$ is $2$-UDA if and only if one of the following conditions holds, (a) $\rho_{A_{n-3}A_{n-2}}$ or $\rho_{A_{n-1}A_{n}}$ is pure, and $\otimes_{i=1}^{n-4}\rho_{A_{i}}\otimes\rho_{A_{j_{1}}A_{j_{2}}}$ satisfies (ii); (b) at least one of $\rho_{A_{j}}$ is pure, for j=n3,,nj=n-3,...,n, and $\rho=\otimes_{i=1}^{n-4}\rho_{A_{i}}\otimes\rho_{A_{j}}\otimes\rho_{A_{l}}\otimes\rho_{A_{m_{1}}A_{m_{2}}}$ satisfies (ii); (c) $\rho_{A_{n-3}A_{n-2}}$ and $\rho_{A_{n-1}A_{n}}$ are rank-two, $\rho_{A_{n-3}A_{n-2}}\otimes\rho_{A_{n-1}A_{n}}$ is LU equivalent to $(H\oplus 0)\otimes(D\oplus 0)$ in (15), and $\otimes_{i=1}^{n-4}\rho_{A_{i}}$ is pure.
The proof of Proposition 23 is given in Appendix D. We shall emphasize that a $k$-UDA state is also $(k+1)$-UDA. This implies that $\rho=\otimes_{i=1}^{n-2}\rho_{A_{i}}\otimes\rho_{A_{n-1}A_{n}}$ is $k$-UDA if the conditions in Proposition 23 (ii) are satisfied.
We turn to consider the UDA property of general $n$-partite states $\rho_{A_{1}A_{2}...A_{n}}$.
Proposition 24
(i) Every full-rank $n$-partite mixed state supported on $\mathbb{C}^{d_{1}}\otimes\mathbb{C}^{d_{2}}\otimes...\otimes\mathbb{C}^{d_{n}}$ is not $k$-UDA, for $1\leq k\leq n-1$.
(ii) Every $n$-qubit mixed state of rank $2^{n}-1$ is not 2-UDA.
Proof.
(i) If $\rho$ is full-rank, then the range ${\cal R}(\rho)$ includes the quasi-GHZ state $a|0\rangle^{\otimes n}+b|1\rangle^{\otimes n}$, which is not $k$-UDA. Thus $\rho$ is not 2-UDA.
(ii) Let $\rho$ be an $n$-qubit mixed state. Suppose $\mathop{\rm rank}\rho=2^{n}-1$. We first prove that $\rho$ is not 2-UDA for $n=3$. We claim that the range of $\rho$ has a pure state $|\gamma\rangle$ LU equivalent to the quasi-GHZ state $a|000\rangle+b|111\rangle$ with $ab\neq 0$. Up to LU equivalence, the kernel of a seven-dimensional space $V\subseteq\mathbb{C}^{8}$ satisfies that $V^{\perp}=\text{span}\{a_{0}|000\rangle+a_{1}|001\rangle+a_{2}|010\rangle+e^{i\theta}a_{3}|011\rangle+a_{4}|111\rangle\}$ acin2000generalize. The state $|\beta\rangle=a_{2}^{*}|001\rangle+\frac{a_{1}^{*}a_{2}}{x^{*}}|110\rangle-a_{1}^{*}|010\rangle+x|101\rangle$ is LU equivalent to the quasi-GHZ state and $|\beta\rangle\perp V^{\perp}$. Therefore, the non-2-UDA state $|\beta\rangle\in V$.
It remains to prove the claim for $n$-qubit states $\rho$ of rank $2^{n}-1$ and $n>3$. It suffices to show that the range of $\rho$ has a pure state LU to the state $|\gamma\rangle|0\rangle^{\otimes(n-3)}$. Let $\ker\rho$ be spanned by the pure state $|\psi\rangle=\sum_{(j_{1},...,j_{n-3})\neq(0,...,0)}|\alpha_{j_{1}...j_{n-3}}\rangle|j_{1}...j_{n-3}\rangle+|\beta\rangle|0\rangle^{\otimes n-3}$. Using the proof for $n=3$, we can assume that $|\beta\rangle$ is orthogonal to $|\gamma\rangle$. Hence $|\psi\rangle$ is orthogonal to $|\gamma\rangle|0\rangle^{\otimes(n-3)}$. We have proven the claim.   $\sqcap\sqcup$
With respect to any unitarily invariant induced measure $\mu_{N,K}$ with $K\geq N$ (in particular the Hilbert–Schmidt or Bures measure), the set of rank-deficient density matrices has measure zero. Indeed, the joint eigenvalue density Zyczkowski2001induced is absolutely continuous on the eigenvalue simplex and vanishes on its boundary. This implies the following fact.
Corollary 25
Almost all $n$-partite mixed states are not $k$-UDA.
It is shown that pure states are typically UDA, which allows local quantum state tomography qin2024state. Corollary 25 shows that the UDA property is unavoidable in the certification of mixed states by such kind of tomography.
The determination of $n$-qubit states that are $k$-UDA is an essential task in quantum state tomography. The full characterization of bipartite to $n$-qubit states we presented in this paper provides a systematic method for determining that some states are not $k$-UDA, as shown in FIG 1. This facilitates narrowing the scope of $k$-UDA states, thereby streamlining their construction and verification. Our goal is determine a $k$-UDA state by known results, or transform a state on more systems and with higher rank into a set of its ”ingredients”, i.e., the marginals of the global state or the elements in its range. Obviously, these states on fewer systems or with lower rank are easier to deal with, as our complete characterization of some multipartite results. The process consists of three steps as follows.
  • STEP 1 (Determination of UDA states) Let $\rho$ be a global state to be determined. By Lemma 5 (i), if $\rho$ is an element of the face of ${\cal R}(\gamma)$ and $\gamma$ is $k$-UDA, then $\rho$ is $k$-UDA. Otherwise, $\rho$ is turned to the non-$k$-UDA verification.
  • STEP 2 (Reduction of systems): All $m$-marginals of $\rho$ are verified by their rank, for $m\leq k$. If the $m$-marginals $\rho_{{\cal S}_{j}}$ are pure, then any state $k$-compatible with $\rho$ is given by $\sigma=\otimes_{j}\rho_{{\cal S}_{j}}\otimes\sigma_{(\cup_{j}{\cal S}_{j})^{c}}$. Then $\rho$ is $k$-UDA iff the remaining part of $\sigma$, i.e. $\sigma_{(\cup_{j}{\cal S}_{j})^{c}}$ is fully determined. That is, the state $\rho_{1}:=\rho_{(\cup_{j}{\cal S}_{j})^{c}}$ on less systems is $k$-UDA. Otherwise, if none of the $m$-partite marginals of $\rho$ is pure, we choose $\rho_{1}=\rho$.
  • STEP 3 (Reduction of rank): $\rho_{1}$ is decomposed to the convex combination of some states $\alpha^{(i)}$, i.e. $\rho=\sum_{i}p_{i}\alpha^{(i)}$. The rank of $\alpha^{(i)}$ will be less than that of $\rho_{i}$ for most cases. The simplification of $\alpha^{(i)}$ is applied by LU equivalence. The set $B$ of states to be verified is constructed as follows. If $\alpha^{(i)}$ is a product state, i.e. $\alpha^{(i)}=\beta^{(i)}_{\cal K}\otimes\beta^{(i)}_{{\cal K}^{c}}$, then $B=\{\alpha^{(i)}\}\cup\{\beta^{(i)}_{\cal K},\beta^{(i)}_{{\cal K}^{c}}\}$; otherwise, $B=\{\alpha^{(i)}\}$. Using Lemma 4 and 5, if there is an element of $B$ that is not $k$-UDA by the necessary and sufficient conditions in Propositions 8 and 23, and Theorems 14, 20 and 22, we conclude that $\rho$ is not $k$-UDA; otherwise, we choose each element $B$ as $\rho$, and repeat STEP 1 to find possible states that is not $k$-UDA.
From this process, one can see that this method can partially conclude or exclude that a state is $k$-UDA. Even if no non-$k$-UDA component is detected through this procedure, the global state $\rho$ may still fail to be $k$-UDA. What can be stated with certainty is that this approach significantly reduces the complexity of assessing whether a multipartite mixed state is $k$-UDA by decomposing a high-rank many-body quantum state into a sequence of lower-rank, smaller subsystem components. Further, this process provides a systematic construction of multipartite states that are not $k$-UDA. This method implies that a state on fewer systems is vital for determining the $k$-UDA property, which is exactly what we have shown in this paper.
Figure 1: Process for the determination of $k$-UDA states. Here $\sigma$ is any state $k$-compatible with $\rho$. The first to third step are shown in the cyan, blue and yellow boxes, respectively.
IV Additivity of mixed UDA states
In this section, we analyze the additivity of bipartite 1-UDA, three-qubit 2-UDA and some $n$-partite states, which is fully characterized in Proposition 8 and Theorem 14. It is shown that the additivity of mixed UDA states can only be satisfied by imposing additional conditions on their marginals. We shall emphasize that UDA additivity of two states is of practical benefits. On the one hand, if the tensor product of two $k$-UDA states are still $k$-UDA, then the complexity for the reconstruction of the resulting states only increases linearly, instead of exponentially. On the other hand, entanglement structure of such states are deduced by their local information since it is certified that the particles of such states cannot be genuinely entangled.
Using Proposition 8, we show the additivity of bipartite 1-UDA states. It is shown that such additivity can be satisfied if and only if one of the states is a pure product state.
Theorem 26
Suppose $\alpha$ and $\beta$ are two bipartite 1-UDA states of systems $(A_{1}A_{2})$ and $(B_{1}B_{2})$, respectively. Then $\alpha\otimes\beta$ is a 1-UDA state of the system $(A_{1}A_{2}B_{1}B_{2})$ if and only if one of the two states $\alpha$ and $\beta$ is a pure product state.
Proof.
Let $\rho_{A_{1}A_{2}B_{1}B_{2}}=\alpha\otimes\beta$. Suppose $\sigma_{A_{1}A_{2}B_{1}B_{2}}$ is a state compatible with $\rho$ by all 1-marginals. From Proposition 8, we have
$\alpha_{A_{1}A_{2}}=|\psi\rangle\!\langle\psi|_{A_{j_{1}}}\otimes\alpha_{A_{j_{2}}},$ (16)
$\beta_{B_{1}B_{2}}=|\eta\rangle\!\langle\eta|_{B_{k_{1}}}\otimes\beta_{B_{k_{2}}},$ (17)
up to system permutations, where $j_{1},j_{2},k_{1},k_{2}\in\{1,2\}$, $j_{1}\neq j_{2}$ and $k_{1}\neq k_{2}$. Then $\sigma_{A_{j_{1}}}=|\psi\rangle\!\langle\psi|$, $\sigma_{A_{j_{2}}}=\alpha_{A_{j_{2}}}$, $\sigma_{B_{k_{1}}}=|\eta\rangle\!\langle\eta|$, $\sigma_{B_{k_{2}}}=\beta_{B_{k_{2}}}$. From Lemma 2, we obtain that
$\sigma_{A_{1}A_{2}B_{1}B_{2}}=|\psi\rangle\!\langle\psi|_{A_{j_{1}}}\otimes|\eta\rangle\!\langle\eta|_{B_{k_{1}}}\otimes\sigma_{A_{j_{2}}B_{k_{2}}}.$ (18)
The state $\sigma=\rho$ if and only if $\sigma_{A_{j_{2}}B_{k_{2}}}=\alpha_{A_{j_{2}}}\otimes\beta_{B_{k_{2}}}=\sigma_{A_{j_{2}}}\otimes\sigma_{B_{k_{2}}}$, i.e. $\sigma_{A_{j_{2}}B_{k_{2}}}$ is a 1-UDA state. Using Proposition 8, $\sigma_{A_{j_{2}}B_{k_{2}}}$ is 1-UDA if and only if one of $\alpha_{A_{j_{2}}}$ and $\beta_{B_{k_{2}}}$ is pure, that is, $\alpha$ or $\beta$ is a pure product state. This completes the proof.   $\sqcap\sqcup$
The additivity of two three-qubit 2-UDA states is rather complicated. By Theorem 14, one can determine whether a product state $\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}$ is 2-UDA directly. We use this fact to investigate the additivity of mixed UDA states. Let $\alpha=\alpha_{A_{1}}\otimes\alpha_{A_{2}A_{3}}$ and $\beta=\beta_{B_{1}}\otimes\beta_{B_{2}B_{3}}$ be two 2-UDA states. We consider the necessary and sufficient condition that $\alpha\otimes\beta$ is 2-UDA. Using Theorem 14, up to local equivalence, $\alpha$ and $\beta$ can be assumed to be in the following form,
$\alpha_{1}=|\psi_{1}\rangle\!\langle\psi_{1}|_{A_{1}}\otimes\alpha_{A_{2}A_{3}},$ (19)
$\alpha_{2}=|\psi_{2}\rangle\!\langle\psi_{2}|_{A_{2}}\otimes\alpha_{A_{1}}\otimes\alpha_{A_{3}}\;$ (20)
$(\text{or }\alpha_{2}=|\psi_{3}\rangle\!\langle\psi_{3}|_{A_{3}}\otimes\alpha_{A_{1}}\otimes\alpha_{A_{2}}),$ (21)
$\alpha_{3}=\alpha_{A_{1}}\otimes|\eta\rangle\!\langle\eta|_{A_{2}A_{3}},$ (22)
$\alpha_{4}=\alpha_{A_{1}}\otimes(H\oplus 0)_{A_{2}A_{3}},\text{ for }\mathop{\rm rank}(\alpha_{A_{1}})=2,$ (23)
where the diagonal elements of $H$ in (15) are positive. Similarly, we have
$\beta_{1}=|\xi_{1}\rangle\!\langle\xi_{1}|_{B_{1}}\otimes\beta_{B_{2}B_{3}},$ (24)
$\beta_{2}=|\xi_{2}\rangle\!\langle\xi_{2}|_{B_{2}}\otimes\beta_{B_{1}}\otimes\alpha_{B_{3}}\;$ (25)
$(\text{or }\beta_{2}=|\xi_{3}\rangle\!\langle\xi_{3}|_{B_{3}}\otimes\beta_{B_{1}}\otimes\alpha_{B_{2}}),$ (26)
$\beta_{3}=\beta_{B_{1}}\otimes|\mu\rangle\!\langle\mu|_{B_{2}B_{3}},$ (27)
$\beta_{4}=\beta_{B_{1}}\otimes(D\oplus 0)_{B_{2}B_{3}},\text{ for }\mathop{\rm rank}(\beta_{B_{1}})=2,$ (28)
with the positive diagonal elements of $D$ in (15). Suppose $\sigma$ is a state compatible with $\alpha_{j}\otimes\beta_{k}$ by all possible 2-marginals. It suffices to investigate the 2-UDA property of $\alpha_{j}\otimes\beta_{k}$. Without loss of generality, this derives ten cases technically, and the following analysis holds, up to swapping the $\alpha_{A_{j}}$’s with the $\beta_{B_{j}}$’s.
  1. 1.
    Suppose $\sigma$ is compatible with $\alpha_{1}\otimes\beta_{1}$. From Lemma 2, we have $\sigma=|\psi_{1}\rangle\!\langle\psi_{1}|_{A_{1}}\otimes|\xi_{1}\rangle\!\langle\xi_{1}|_{B_{1}}\otimes\sigma_{A_{2}A_{3}B_{2}B_{3}}$, where $\sigma_{A_{2}A_{3}}=\alpha_{A_{2}A_{3}}$, $\sigma_{B_{2}B_{3}}=\beta_{B_{2}B_{3}}$ and $\sigma_{A_{m}A_{n}}=\alpha_{A_{m}}\otimes\beta_{B_{n}}$. One can verify that $\alpha_{1}\otimes\beta_{1}$ is 2-UDA if and only if $\sigma_{A_{2}A_{3}B_{2}B_{3}}=\alpha_{A_{2}A_{3}}\otimes\beta_{B_{2}B_{3}}$. That is, $\alpha_{A_{2}A_{3}}\otimes\beta_{B_{2}B_{3}}$ is 2-UDA and thus satisfies the conditions in Theorem 20.
    Similarly, we consider the remaining nine cases and obtain the following facts.
  2. 2.
    $\alpha_{1}\otimes\beta_{2}$ is 2-UDA if and only if $\alpha_{A_{2}A_{3}}\otimes\beta_{B_{1}}\otimes\beta_{B_{3}}$ is 2-UDA. This is analyzed in Proposition 17.
  3. 3.
    $\alpha_{1}\otimes\beta_{3}$ is 2-UDA if and only if $\beta_{B_{1}}\otimes\alpha_{A_{2}A_{3}}$ is 2-UDA. That is, the resulting state $\beta_{B_{1}}\otimes\alpha_{A_{2}A_{3}}$ satisfies one of the conditions in Theorem 14.
  4. 4.
    $\alpha_{1}\otimes\beta_{4}$ is 2-UDA if and only if $\alpha_{A_{2}A_{3}}\otimes\beta_{B_{1}}\otimes(D\oplus 0)_{B_{2}B_{3}}$ is 2-UDA, which is analyzed in Theorem 22.
  5. 5.
    $\alpha_{2}\otimes\beta_{2}$ is 2-UDA if and only if $\alpha_{A_{1}}\otimes\alpha_{A_{3}}\otimes\beta_{B_{1}}\otimes\beta_{B_{3}}$ is 2-UDA. This is shown in Proposition 15.
  6. 6.
    $\alpha_{2}\otimes\beta_{3}$ is 2-UDA if and only if $\alpha_{A_{1}}\otimes\alpha_{A_{3}}\otimes\beta_{B_{1}}$ is 2-UDA. Using Proposition 9, $\alpha_{2}\otimes\beta_{3}$ is 2-UDA iff one of $\alpha_{A_{1}}$, $\alpha_{A_{3}}$ and $\beta_{B_{1}}$ is pure.
  7. 7.
    $\alpha_{2}\otimes\beta_{4}$ is 2-UDA iff $\alpha_{A_{1}}\otimes\alpha_{A_{3}}\otimes\beta_{B_{1}}\otimes(D\oplus 0)_{B_{2}B_{3}}$ is 2-UDA. This is considered in Proposition 23.
  8. 8.
    $\alpha_{3}\otimes\beta_{3}$ is 2-UDA without additional conditions.
  9. 9.
    $\alpha_{3}\otimes\beta_{4}$ is 2-UDA if and only if $\alpha_{A_{1}}\otimes\beta_{B_{1}}\otimes(D\oplus 0)_{B_{2}B_{3}}$ is 2-UDA by Theorem 20.
  10. 10.
    $\alpha_{4}\otimes\beta_{4}$ is 2-UDA iff $\alpha_{A_{1}}\otimes(H\oplus 0)_{A_{2}A_{3}}\otimes\beta_{B_{1}}\otimes(D\oplus 0)_{B_{2}B_{3}}$ is 2-UDA. From Theorem 22, this state is 2-UDA iff both $\alpha_{A_{1}}$ and $\beta_{B_{1}}$ are pure.
By the preceding analysis, we obtain the necessary and sufficient condition under which the additivity of two three-qubit UDA states holds.
Theorem 27
Suppose $\alpha$ and $\beta$ are two three-qubit 2-UDA states of systems $(A_{1}A_{2}A_{3})$ and $(B_{1}B_{2}B_{3})$, respectively. Then $\alpha\otimes\beta$ is a 2-UDA state of the system $(A_{1}A_{2}A_{3}B_{1}B_{2}B_{3})$ if and only if one of the following conditions holds, up to swapping $\alpha_{A_{m}}$ with $\beta_{B_{m}}$,
(i) $\alpha_{A_{2}A_{3}}$ and $\beta_{B_{2}B_{3}}$ are pure;
(ii) $\alpha_{A_{j_{1}}}$ and $\beta_{B_{2}B_{3}}$ (resp. $\alpha_{A_{j_{1}}}$ and $\beta_{B_{k_{1}}}$) are pure, and the state $\alpha_{A_{j_{2}}A_{j_{3}}}\otimes\beta_{B_{1}}$ (resp. $\alpha_{A_{j_{2}}A_{j_{3}}}\otimes\beta_{B_{k_{2}}B_{k_{3}}}$) is 2-UDA, satisfying Theorem 14 (resp. Theorem 20);
(iii) $\alpha_{A_{2}A_{3}}$ (resp. $\alpha_{A_{j_{1}}}$) is pure, $\mathop{\rm rank}(\beta_{B_{2}B_{3}})=2$ and is LU equivalent to $D\oplus 0$ in (15). The state $\alpha_{A_{1}}\otimes\beta_{B_{1}}\otimes(D\oplus 0)$ (resp. $\alpha_{A_{j_{2}}A_{j_{3}}}\otimes\beta_{B_{k_{2}}}\otimes(D\oplus 0)$) is 2-UDA satisfying Theorem 20 (resp. Theorem 22).
(iv) $\alpha_{A_{1}}$ and $\beta_{B_{1}}$ are pure, $\mathop{\rm rank}(\alpha_{A_{2}A_{3}})=\mathop{\rm rank}(\beta_{B_{2}B_{3}})=2$, and $\alpha_{A_{2}A_{3}}\otimes\beta_{B_{2}B_{3}}$ is LU equivalent to $(H\oplus 0)\otimes(D\oplus 0)$ in (15).
We recall that Theorem 1 in add2023shen shows that if two mixed states $\alpha$ and $\beta$ are both $k$-UDA and one of them is pure, then $\alpha\otimes\beta$ is $k$-UDA directly. Obviously, this fact is included in Theorem 27 (ii), where a pure $\alpha_{A_{j_{2}}A_{j_{3}}}\otimes\beta_{B_{1}}$ naturally satisfies the conditions in Theorem 14. It is also shown that the purity of $\alpha$ or $\beta$ is a rather strict condition for mixed states since the additivity also holds for mixed UDA $\alpha$ and $\beta$. Theorem 27 implies that the rank of $\alpha$ and $\beta$ is less than four and the rank of $\alpha\otimes\beta$ is less than 4 if their UDA additivity holds. Note that $\alpha\otimes\beta$ is a six-qubit state whose rank is at most 64. Hence, UDA additivity holds only for states with relatively low rank.
We shall consider the UDA additivity of some $n$-partite states by Proposition 23. We assume that the $n$-partite states $\alpha_{1}=\otimes_{i=1}^{n}\alpha_{A_{i}}$ and $\beta_{1}=\otimes_{j=1}^{m}\beta_{B_{j}}$ be two $k$-UDA states, then at least nkn-k and mkm-k 1-marginals of $\alpha_{1}$ and $\beta_{1}$ are pure, respectively. Proposition 23 (i) further shows that for $k_{1}\geq n-k$, $k_{2}\geq m-k$ and $k_{1}+k_{2}\geq m+n-k$, $\alpha_{1}\otimes\beta_{1}$ is $k$-UDA iff at least $k_{1}$ and $k_{2}$ 1-marginals of $\alpha_{1}$ and $\beta_{1}$ are pure, respectively. Similarly, the UDA additivity of the following mixed states can be derived. The additivity of $\otimes_{i=1}^{n}\alpha_{A_{i}}$ and $\otimes_{j=1}^{m-2}\beta_{B_{j}}\otimes\beta_{B_{m-1}B_{m}}$ (resp. $\otimes_{j=1}^{m-4}\beta_{B_{j}}\otimes\beta_{B_{m-3}B_{m-2}}\otimes\beta_{B_{m-1}B_{m}}$) can be derived directly by Proposition 23 (ii)( resp. (iii)). Besides, the UDA additivity of $\otimes_{i=1}^{n-2}\alpha_{A_{i}}\otimes\alpha_{A_{n-1}A_{n}}$ and $\otimes_{j=1}^{m-2}\beta_{B_{j}}\otimes\beta_{B_{m-1}B_{m}}$ is also shown by Proposition 23 (iii). The preceding analysis shows that the UDA additivity can not hold without additional conditions. The necessary and sufficient conditions of $\alpha$ and $\beta$, which derive the additivity hold is given above. This answers the question we proposed at the beginning of this work. Obviously, these conditions naturally require the rank of two states satisfying UDA additivity much lower.
V Applications
In this section, we show the physical applications of mixed UDA states and their characterization. First, we show that the determination lowers the complexity of mixed-state tomography. Next, we introduce the applications from the perspectives of entanglement, quantum communication, error correction, and many-body quantum phases.
An $n$-qubit density matrix $\rho$ is parameterized by
$\rho=\frac{1}{2^{n}}\sum_{i_{1},\ldots,i_{n}=0}^{3}\pi_{i_{1}\cdots i_{n}}\,\sigma_{i_{1}}^{(1)}\otimes\cdots\otimes\sigma_{i_{n}}^{(n)}.$
The reconstruction of $\rho$ usually requires the expectation values $\pi_{i_{1}\cdots i_{n}}$ of $D_{U}=4^{n}-1$ Pauli observables and many samples of $\rho$. The complexity is later lowered in fully quantum state tomography (FQST), where $D_{F}=3^{n}$ global Pauli observables are required by ignoring $\sigma_{0}$ in each qubit. However, FQST is unfeasible for large systems. By using the complete characterization of UDA states, global states can be reconstructed from measurements on $k$-marginals, a technique known as local measurements on reduced density matrices (LQST). Generally, LQST requires measuring $D_{L}=3^{k}\binom{n}{k}$ local observables. The Parallel quantum state tomography (PQST) protocol for $n$-qubit states experiment2024hu exploits parallel Pauli measurements to efficiently obtain all $k$-marginals, and reconstructs the global state via tensor-network learning. By choosing $k=2$, PQST requires $D_{P}^{(2)}=3+6\lceil\log_{2}n\rceil$ observables.
LQST is grounded in the fact that pure states are generally UDA. However, almost all mixed states are not UDA by Corollary 25, and the complete characterization of mixed UDA states is necessary in their LQST. Besides, PQST reconstructs the global state via tensor-network learning, without assuming the UDA property. Using the characterization of UDA states, high-fidelity-PQST will be realized with fewer observables. In fact, the set of global Pauli observables ${\cal M}_{\text{PQST}}$ is sufficient to reconstruct $k$-marginals, which implies the observables used in PQST of mixed UDA states are less than generic states, i.e. $D_{\text{PQST}}^{\text{UDA}}\leq D_{\text{PQST}}^{\text{generic}}$. Further, a UDA state removes unidentifiable directions in the reconstruction. Statistical noise cannot induce large deviations, thereby reducing estimation variance. This improves statistical efficiency and leads to higher reconstruction fidelity. Since the measurement cost and noisy error grow with $k$, we consider the tomography of 2-UDA states in four types of tomography, and compare the number of observables in TABLE 1. One can see that the number of observables increases exponentially for usual tomography and FQST. Using the characterization of mixed UDA state, LQST requires much fewer observables, which grows polynomially with $n$. Combining with the results of mixed UDA states, the observables in PQST may be less than $D_{P}$ in TABLE 1. This significantly reduces the measurement and computational resources required for state reconstruction, especially in large systems.
2-UDA state $D_{U}$ $D_{F}$ $D_{L}$ $D_{P}$
Three-qubit state $\rho_{A_{1}A_{2}A_{3}}$ in Theorem 20 63 27 27 1515
Four-qubit state $\rho_{A_{1}A_{2}}\otimes\rho_{A_{3}A_{4}}$ in Theorem 20 255 81 54 1515
Five-qubit state $\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}}$ in Theorem 22 1023 243 90 2121
Six-qubit state $\alpha\otimes\beta$ in Theorem 27 4095 729 135 2121
$n$-qubit state $\otimes_{i=1}^{n-4}\rho_{A_{i}}\otimes\rho_{A_{n-3}A_{n-2}}\otimes\rho_{A_{n-1}A_{n}}$ in Proposition 23 $4^{n}-1$ $3^{n}$ $3^{k}\binom{n}{k}$ $3+6\lceil\log_{2}n\rceil$
Table 1: The number of observables required in four types of quantum tomography. Here $D_{U}$, $D_{F}$, $D_{L}$, $D_{P}$ denote the number of observables used for usual tomography, FQST, LQST, PQST, respectively.
In addition to QST, the characterization of mixed UDA states sheds light on other applications as follows.
  • Entanglement detection and distribution. When an entangled state is UDA, entanglement can be certified indirectly through local measurements yu2020multi; shi2025entangle. Otherwise, a non-$k$-UDA state implies the irreducible $(k+1)$-partite correlations irr2008zhou.
  • Security in quantum communication and cryptography. An adversary may attempt to distribute a counterfeit global state that reproduces the same local statistics as a desired resource state. If the target state is UDA with respect to the monitored data, such forgery is impossible huang2025certify; swingle2014reconstruct.
  • Quantum error correction and recoverability. A UDA state reflects a form of information redundancy: the global quantum state is uniquely encoded in partial subsystems. This is related to quantum error correction, where recovering a state from partial information is a key requirement gupta2024encode.
  • Characterization of many-body quantum phases. States exhibiting topological order or long-range entanglement are generally not UDA with respect to local reduced states lee2025topo, whereas short-range entangled phases often admit UDA properties. This connects UDA states to the study of quantum phases and phase classification will2025probing.
VI Conclusion
In summary, we have characterized some mixed UDA states, and analyzed their additivity by giving the necessary and sufficient conditions. We have shown the general properties satisfied by a mixed UDA state. Using these facts, we show the complete characterization of mixed product states of bipartite to $n$-partite. A systematic method has been developed for determining $k$-UDA states. It derives the construction of multipartite states that is not $k$-UDA. The UDA additivity of bipartite 1-UDA, three-qubit 2-UDA and some $n$-partite states is studied using the characterization of mixed UDA states. Further, we present the applications of mixed UDA states and their characterization.
Many problems arising from this work can be further explored. First, we know that a full-rank state is not UDA, while a low-rank state may be UDA. It is interesting to analyze the threshold of the rank of a mixed state, where states whose rank exceeds this threshold are not UDA. Second, the necessary and sufficient condition by which arbitrary multipartite states can be further considered. The techniques of purification of mixed states and range inclusion may be utilized. Then the UDA additivity of arbitrary multipartite mixed states can be further investigated.
ACKNOWLEDGMENTS
This work is supported by the NNSF of China (Grant No. 12471427) and the NUS research grant A-8003570-00-00.
Appendix A Proof of preliminary lemmas
Lemma 4 shows that $\alpha\otimes\beta$ is not $k$-UDA if either of $\alpha$ and $\beta$ is not $k$-UDA. Its proof is shown as follows.
Proof.
We denote by $A_{[m]}$ and $B_{[n]}$ the $m$-partite system $(A_{1},...,A_{m})$ and $n$-partite system $(B_{1},...,B_{n})$, respectively. Let $\rho_{A[m]B[n]}=\alpha_{A_{[m]}}\otimes\beta_{B_{[n]}}$. If $\rho$ is $k$-UDA, then for any state $\sigma$ compatible with $\rho$, i.e. $\rho_{{\cal S}}=\sigma_{{\cal S}}$ for ${\cal S}\subset A_{[m]}B_{[n]}$ and $|{\cal S}|=k$, it holds that $\rho=\sigma$. Since $\alpha$ and $\beta$ are both not $k$-UDA, there are states $\eta_{A_{[m]}}$ and $\xi_{B_{[n]}}$ such that $\eta_{{\cal S}_{a}}=\alpha_{{\cal S}_{a}}$ for ${\cal S}_{a}\subset A_{[m]}$ and $\xi_{{\cal S}_{b}}=\beta_{{\cal S}_{b}}$ for ${\cal S}_{b}\subset B_{[m]}$, and $\alpha\neq\eta$, $\beta\neq\xi$. We claim that $\eta\otimes\xi$ has the same $k$-marginals as that of $\alpha\otimes\beta$. If ${\cal S}\subset A_{[m]}$ or ${\cal S}\subset B_{[n]}$, the claim holds by choosing ${\cal S}={\cal S}_{a}$ or ${\cal S}={\cal S}_{b}$. Otherwise, suppose ${\cal S}=(S_{1},S_{2})$ and $S_{1}\in A_{[m]}$, $S_{2}\in B_{[n]}$. Let ${\cal G}_{1}=\{S_{1},S_{l}\}\subset A_{[m]}$ and ${\cal G}_{2}=\{S_{2},S_{j}\}\subset B_{[n]}$. Then
$\hskip-8.5359pt(\eta\otimes\xi)_{{\cal S}}=$ $\mathop{\rm Tr}_{\{A_{[m]}\backslash S_{1}\}\cup\{B_{[n]}\backslash S_{2}\}}(\eta\otimes\xi)$
$=$ $\mathop{\rm Tr}_{S_{k}}\big[\mathop{\rm Tr}_{\{A_{[m]}\backslash{\cal G}_{1}\}}(\eta)\big]\otimes\mathop{\rm Tr}_{S_{j}}\big[\mathop{\rm Tr}_{\{B_{[n]}\backslash{\cal G}_{2}\}}(\xi)\big]$
$=$ $\mathop{\rm Tr}_{S_{k}}(\alpha_{{\cal G}_{1}})\otimes\mathop{\rm Tr}_{S_{j}}(\beta_{{\cal G}_{2}})=\alpha_{S_{1}}\otimes\beta_{S_{2}}$
$=$ $(\alpha\otimes\beta)_{{\cal S}}=\rho_{\cal S}.$
We choose $\sigma=\eta\otimes\xi$. Then $\sigma_{{\cal S}}=\rho_{{\cal S}}$ for any ${\cal S}\subset A_{[m]}B_{[n]}$. Since $\rho$ is $k$-UDA, we have $\rho=\sigma$. However, $\alpha\neq\eta$ and $\beta\neq\xi$ implies that $\alpha\otimes\beta\neq\eta\otimes\xi$. In fact, we assume that $\alpha=(a_{ij})$ and $\beta=(b_{ij})$ be two normalized quantum states. From $\alpha\neq\beta$, there is $a_{ij}\neq b_{ij}$ and without loss of generality, $a_{ij}\neq 0$. Suppose $b_{ij}=ha_{ij}$ for $h\neq 1$. If $\alpha\otimes\beta=\eta\otimes\xi$, then $a_{ij}\beta=b_{ij}\xi$. So $\beta=h\xi$ and $\beta$ is not normalized. Hence we have $\rho\neq\sigma$. It is a contradiction.   $\sqcap\sqcup$
Lemma 5 shows three facts. First, a state whose range includes a non-$k$-UDA state is itself not $k$-UDA. Second, two states that have the same range are $k$-UDA or not at the same time. Third, if two states are $k$-compatible, then their marginals are also $k$-compatible. The proof is given below.
Proof.
(i) Suppose $\rho$ and $\sigma$ are on ${\cal H}_{A_{1}A_{2}...A_{n}}$. Since $\rho$ is not $k$-UDA, there is a state $\rho^{\prime}$ such that $\rho_{\cal S}=\rho^{\prime}_{{\cal S}}$ for all ${\cal S}\subset(A_{1}A_{2}...A_{n})$ and $|{\cal S}|=k$. Then $\lambda\rho_{\cal S}+(1-\lambda)\sigma_{\cal S}=\lambda\rho^{\prime}_{\cal S}+(1-\lambda)\sigma_{\cal S}$ implies that $\lambda\rho^{\prime}+(1-\lambda)\sigma$ is $k$-compatible with $\lambda\rho+(1-\lambda)\sigma$, and they are different states.
(ii) From ${\cal R}(\beta)\subset{\cal R}(\alpha)$, we have $\alpha=(\alpha-\epsilon\beta)+\epsilon\beta$ for $\epsilon>0$ small enough. Using Lemma 3, we have $\alpha-\epsilon\beta\geq 0$ is a (unnormalized) state. From Lemma 5, if $\beta$ is not 2-UDA, then $\alpha$ is not 2-UDA. Similarly, by ${\cal R}(\alpha)\subset{\cal R}(\beta)$, $\alpha$ is not 2-UDA implies that $\beta$ is not 2-UDA.
(iii) Since $\rho$ and $\sigma$ are $k$-compatible, we can trace out the systems ${\cal S}^{c}$ on $\rho$ and $\sigma$, and the resulting states are also $k$-compatible. This completes the proof.   $\sqcap\sqcup$
Lemma 6 shows that a $m$-partite state $\rho=\rho_{A_{1}A_{2}...A_{n}}\otimes\rho_{A_{n+1}A_{n+2}...A_{m}}$ satisfies ${\cal R}(\sigma)\subseteq{\cal R}(\rho)$ for any $\sigma$ that is $n$-compatible with $\rho$. The proof is present here.
Proof.
Suppose ${\cal R}(\rho_{A_{1}A_{2}...A_{n}})=\text{span}\{|\alpha_{k}\rangle:k=1,2,...,r\}$. Let $|\psi\rangle$ be the purification of $\rho_{A_{1}A_{2}...A_{n}}$ on systems $A_{1}A_{2}...A_{m}$. Then
$|\psi\rangle=\sum_{k}|\alpha_{k}\rangle_{A_{1}...A_{n}}|\beta_{k}\rangle_{A_{n+1}...A_{m}}.$
For any $n$-partite channel $\Lambda(\cdot)=\sum_{j=1}^{k}M_{j}(\cdot)M_{j}^{\dagger}$ acting on ${\cal H}_{A_{n+1}...A_{m}}$, we have
$\sigma=$ $\mathbb{I}_{A_{1}A_{2}...A_{n}}\otimes\Lambda_{A_{n+1}A_{n+2}...A_{2n}}(|\psi\rangle\!\langle\psi|)$ (29)
$=$ $\sum_{j}(\mathbb{I}\otimes M_{j})|\psi\rangle\!\langle\psi|(\mathbb{I}\otimes M_{j}^{\dagger}).$
Since $\rho_{A_{n+1}...A_{m}}$ is $n$-UDA, we have $\sigma_{A_{n+1}...A_{m}}=\rho_{A_{n+1}...A_{m}}$. Hence, $M_{j}|\beta_{k}\rangle\in{\cal R}(\rho_{A_{n+1}...A_{m}})$. Hence ${\cal R}(\sigma)\subseteq{\cal R}(\rho)$.   $\sqcap\sqcup$
Lemma 7 shows that two states $\alpha\otimes\beta$ and $\beta\otimes\alpha$ are the same time $k$-UDA or not. Its proof is given here.
Proof.
Let $\alpha$ and $\beta$ be two states of systems $(A_{1}A_{2}...A_{m})$ and $(B_{1}B_{2}...B_{n})$, respectively. Suppose $\alpha\otimes\beta$ is not 2-UDA. There is a different $(m+n)$-partite state $\sigma_{A_{[m]}B_{[n]}}$ that is $k$-compatible with $\alpha\otimes\beta$, i.e. $\sigma_{A_{\cal K}B_{\cal P}}=\alpha_{A_{\cal K}}\otimes\beta_{B_{\cal P}}$ with $|A_{\cal K}|+|B_{\cal P}|=k$. Obviously, $\sigma_{B_{\cal P}A_{\cal K}}=\beta_{B_{\cal P}}\otimes\alpha_{A_{\cal K}}$. That is, the $k$-partite marginals of the state $\sigma_{B_{[n]}A_{[m]}}$ is the same as that of $\beta\otimes\alpha$. Since $\sigma_{A_{[m]}B_{[n]}}\neq\alpha\otimes\beta$, we have $\sigma_{B_{[n]}A_{[m]}}\neq\beta\otimes\alpha$. This completes the proof.   $\sqcap\sqcup$
Appendix B Proof of tripartite UDA states
In Proposition 9, we show that $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}$ with diagonal $\rho_{A_{2}A_{3}}$ is 2-UDA if and only if one of the six 1- and 2-marginals of $\rho$ is pure. We show the proof of Proposition 9 as follows.
Proof.
The if part is trivial by Lemma 2. It remains to show the ”only if” part. Suppose the tripartite state $\rho$ is supported on $\mathbb{C}^{d_{1}}\otimes\mathbb{C}^{d_{2}}\otimes\mathbb{C}^{d_{3}}$. Since the UDA property is invariant up to local unitaries, we assume that
$\rho:=$ $\mathop{\rm diag}\{p_{1},p_{2},...,p_{d_{1}}\}_{A_{1}}\otimes\rho_{A_{2}A_{3}}$ (30)
$=$ $\mathop{\rm diag}\{p_{1},p_{2},...,p_{d_{1}}\}_{A_{1}}\otimes\mathop{\rm diag}\{M_{1},M_{2},\cdots,M_{d_{2}}\},$
where $p_{j}\in[0,1]$ and $M_{k}:=\mathop{\rm diag}\{q_{(k-1)d_{3}+1},...,q_{kd_{3}}\}$. If none of the 1- and 2-marginals of $\rho$ is pure, then
$\hskip-17.07182pt\mathop{\rm rank}(\rho_{A_{1}})=\mathop{\rm rank}(\mathop{\rm diag}\{p_{1},p_{2},...,p_{d_{1}}\})>1,$ (31)
$\hskip-17.07182pt\mathop{\rm rank}(\rho_{A_{2}})=\mathop{\rm rank}(\mathop{\rm diag}\{\mathop{\rm Tr}(M_{1}),...,\mathop{\rm Tr}(M_{d_{2}})\})>1,$ (32)
$\hskip-17.07182pt\mathop{\rm rank}(\rho_{A_{3}})=\mathop{\rm rank}(\sum_{k=1}^{d_{2}}M_{k})>1.$ (33)
since all the preceding 1-marginals of $\rho$ are not pure implies that all the 2-marginals $\rho_{A_{1}A_{2}}$, $\rho_{A_{1}A_{3}}$, $\rho_{A_{2}A_{3}}$ are not pure.
From (31)-(33), at least two of $p_{j}$’s and $q_{k}$’s are strictly larger than zero. Suppose $p_{j_{1}},\;p_{j_{2}}>0$ for $1\leq j_{1}<j_{2}\leq d_{1}$ and $q_{k_{1}},\;q_{k_{2}}>0$ with $1\leq k_{1}<k_{2}\leq d_{2}d_{3}$. Then we have a different state $\sigma$ compatible with $\rho$, that is,
$\sigma=$ $\oplus_{l=1}^{j_{1}-1}p_{l}\rho_{A_{2}A_{3}}\oplus T_{1}\oplus_{l=j_{1}+1}^{j_{2}-1}p_{l}\rho_{A_{2}A_{3}}$
$\oplus T_{2}\oplus_{l=j_{2}+1}^{d_{1}}p_{l}\rho_{A_{2}A_{3}},$
where the diagonal elements of $T_{l}$ and are the in turns $p_{j_{l}}q_{1}$,…, $p_{j_{l}}q_{k_{1}}$, …, $p_{j_{l}}q_{k_{2}}$, …, $p_{j_{l}}q_{d_{2}d_{3}}$. The only nonzero off-diagonal elements of $T_{1}$ and $T_{2}$ are at $k_{1}$-th row and $k_{2}$-th column (and also $k_{2}$-th row and $k_{1}$-th column by Hermitian property), it is $x$ and x-x, respectively. For the nonzero element $x$, it satisfies that
$0<|x|\leq\mathop{\rm min}\{p_{j_{1}}\sqrt{q_{k_{1}}q_{k_{2}}},\;p_{j_{2}}\sqrt{q_{k_{1}}q_{k_{2}}}\}$ (35)
The codition (35) derives that $\sigma\geq 0$ and $\sigma\neq\rho$. Hence $\rho$ is not 2-UDA. This completes the proof.   $\sqcap\sqcup$
Proposition 12 fully characterizes whether a state $\rho=\rho_{A_{1}}\otimes(D\oplus 0)$ in (7) is 2-UDA. The proof of Proposition 12 is shown.
Proof.
(i) If $\mathop{\rm rank}(\rho_{A_{1}})=1$ or $q_{2}=0$ or $q_{3}=0$, then one of a single-qubit marginals of $\rho$ is pure. Hence $\rho$ is 2-UDA.
(ii) Up to local unitaries, we assume that $\rho_{A_{1}}=p|0\rangle\!\langle 0|+(1-p)|1\rangle\!\langle 1|$. We show the ”if” part. Let $\sigma$ be the state compatible with $\rho$. Since $\sigma$ equal to all 2-marginals of $\rho$, we obtain that
$\sigma=\begin{bmatrix}pq_{1}&py&pz&0&0&0&0&0\\ py^{*}&pq_{2}&s_{23}&0&0&0&s_{27}&0\\ pz^{*}&s_{23}^{*}&pq_{3}&0&0&s_{36}&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&(1-p)q_{1}&(1-p)y&(1-p)z&0\\ 0&0&s_{36}^{*}&0&(1-p)y^{*}&(1-p)q_{2}&x-s_{23}&0\\ 0&s_{27}^{*}&0&0&(1-p)z^{*}&x^{*}-s_{23}^{*}&(1-p)q_{3}&0\\ 0&0&0&0&0&0&0&0\end{bmatrix}$. (36)
From $|D|=0$, there is a nonzero vector $v=[a,b,c]^{\dagger}$ such that $Dv=0$. That is,
$\begin{cases}q_{1}a+yb+zc=0\\ y^{*}a+q_{2}b+xc=0\\ z^{*}a+x^{*}b+q_{3}c=0.\end{cases}$ (37)
Suppose $B_{1}$ and $B_{2}$ are the three-order principle minors formed from rows and columns (1, 2, 3) and (5, 6, 7) of the $\sigma$, respectively. Note that $B_{1}+B_{2}=D$. We have $B_{1}v+B_{2}v=0$, and hence $v^{*}B_{1}v+v^{*}B_{2}v=0$. From $B_{1},B_{2}\geq 0$, we have $v^{*}B_{1}v\geq 0$ and $v^{*}B_{2}v\geq 0$. Then $v^{*}B_{1}v=v^{*}B_{2}v=0$, and thus $B_{1}v=B_{2}v=0$. Using (37), we have
$B_{1}v=-B_{2}v=[0,c(s_{23}-px),b(s_{23}^{*}-px^{*})]^{\dagger}=0.$ (38)
From the first equality of (37), $b=c=0$ implies that $a=0$ by $q_{1}>0$, and thus $v=0$, which is a contradiction. It holds that $c\neq 0$ or $b\neq 0$ in $v$. By (38), we have $s_{23}=px$.
Next we show that $s_{36}=s_{27}=0$. The positive semidefiniteness of $\sigma$ implies that $|x|\leq\sqrt{q_{2}q_{3}}$, $|y|\leq\sqrt{q_{1}q_{2}}$ and $|z|\leq\sqrt{q_{1}q_{3}}$. This derives four cases, depending on the value of $|x|,|y|$ and $|z|$.
Case 1: $|x|<\sqrt{q_{2}q_{3}}$, $|y|<\sqrt{q_{1}q_{2}}$ and $|z|<\sqrt{q_{1}q_{3}}$ hold simultaneously. From
$C_{1}:=\begin{bmatrix}pq_{1}&py&pz&0\\ py^{*}&pq_{2}&px&s_{27}\\ pz^{*}&px^{*}&pq_{3}&0\\ 0&s_{27}^{*}&0&(1-p)q_{3}\end{bmatrix}\geq 0,$ (39)
$C_{2}:=\begin{bmatrix}pq_{1}&py&pz&0\\ py^{*}&pq_{2}&px&0\\ pz^{*}&px^{*}&pq_{3}&s_{36}\\ 0&0&s_{36}^{*}&(1-p)q_{2}\end{bmatrix}\geq 0,$ (40)
we have $(1-p)p^{3}q_{3}|D|-|s_{27}|^{2}p^{2}(q_{1}q_{3}-|z|^{2})\geq 0$, $(1-p)p^{3}q_{2}|D|-|s_{36}|^{2}p^{2}(q_{1}q_{2}-|y|^{2})\geq 0$. By $|D|=0$, $|y|<\sqrt{q_{1}q_{2}}$ and $|z|<\sqrt{q_{1}q_{3}}$, we obtain that $s_{27}=s_{36}=0$.
Case 2: $|x|=\sqrt{q_{2}q_{3}}$. By the positive semi-definiteness of $\sigma$, the following matrices are positive semidefinite
$\begin{bmatrix}pq_{2}&px&s_{27}\\ px^{*}&pq_{3}&0\\ s_{27}^{*}&0&(1-p)q_{2}\end{bmatrix},\begin{bmatrix}pq_{2}&px&0\\ px^{*}&pq_{3}&s_{36}\\ 0&s_{36}^{*}&(1-p)q_{2}\end{bmatrix}$
We have $p^{2}(1-p)q_{2}(q_{2}q_{3}-|x|^{2})-pq_{3}|s_{27}|^{2}=-pq_{3}|s_{27}|^{2}\geq 0$, and $s_{27}=0$ by $pq_{3}>0$. Similarly, $D_{2}\geq 0$ derives that $s_{36}=0$.
Case 3: $|y|=\sqrt{q_{1}q_{2}}$. We have $s_{27}=s_{36}=0$ by the positivity of the following matrices,
$\begin{bmatrix}pq_{1}&py&0\\ py^{*}&pq_{2}&s_{27}\\ 0&s_{27}^{*}&(1-p)q_{3}\end{bmatrix},\begin{bmatrix}pq_{3}&0&s_{36}\\ 0&(1-p)q_{1}&(1-p)y\\ s_{36}^{*}&(1-p)y^{*}&(1-p)q_{2}\end{bmatrix}$.
Case 4: $|z|=\sqrt{q_{1}q_{3}}$. Then $s_{27}=s_{36}=0$ by the positivity of the following matrices,
$\begin{bmatrix}pq_{1}&pz&0\\ pz^{*}&pq_{3}&s_{36}\\ 0&s_{36}^{*}&(1-p)q_{2}\end{bmatrix},\begin{bmatrix}pq_{2}&0&s_{27}\\ 0&(1-p)q_{1}&(1-p)z\\ s_{27}^{*}&(1-p)z^{*}&(1-p)q_{3}\end{bmatrix}$.
To conclude, we show that $s_{23}=px$ and $s_{27}=s_{36}=0$, i.e. $\sigma=\rho$, by assuming that $|D|=0$. We have proven the ”if” part.
We show the ”only if” part. Suppose $|D|>0$. Then $|x|<\sqrt{q_{2}q_{3}}$, $|y|<\sqrt{q_{1}q_{2}}$ and $|z|<\sqrt{q_{1}q_{3}}$ hold simultaneously. Let $h_{1}=q_{2}q_{3}-|x|^{2}$, $h_{2}=q_{1}q_{2}-|y|^{2}$ and $h_{3}=q_{1}q_{3}-|z|^{2}$. Then $h_{k}>0$. We show that there is another state $\sigma_{1}$ compatible with $\rho$ by all possible 2-marginals. In detail, we set all elements of $\sigma_{1}$ the same as $\rho$, except for $s_{27}$. Now we show that the element $s_{27}$ of $\sigma_{1}$ is not necessarily equal to zero. We check all principal minors of $\sigma_{1}$ that contains the element $s_{27}$, and obtain that
$0<|s_{27}|^{2}\leq f,$ (41)
where $f=\mathop{\rm min}\{(1-p)ph_{2}h_{3}/q_{1}^{2},(1-p)ph_{1}h_{2}/q_{1}q_{2},p(1-p)h_{1}h_{3}/q_{1}q_{3},p(1-p)h_{1}^{2}/q_{2}q_{3},p(1-p)h_{1}t/q_{3}h_{2},p(1-p)t/q_{1},(1-p)ph_{1}t/q_{2}h_{3},p(1-p)t^{2}/h_{2}h_{3}\}$ with $t=|D|>0$. Note that the RHS of (41) is strictly larger than zero. Hence we can choose $s_{27}>0$ satisfying (41) in $\sigma_{1}$. One can verify that $\sigma_{1}$ is compatible with $\rho$ by all possible 2-marginals and $\sigma_{1}\neq\rho$. Hence $\rho$ is not 2-UDA. This completes the proof.
(iii) We show the ”if” part. Let $\sigma$ be the state compatible with $\rho$. Since $\sigma$ equal to all 2-marginals of $\rho$, we obtain that
$\sigma=0\oplus\begin{bmatrix}pq_{2}&s_{23}&0&0&0&s_{27}\\ s_{23}^{*}&pq_{3}&0&0&s_{36}&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&s_{36}^{*}&0&0&(1-p)q_{2}&x-s_{23}\\ s_{27}^{*}&0&0&0&x^{*}-s_{23}^{*}&(1-p)q_{3}\end{bmatrix}\oplus 0.$ (42)
Let
$D_{1}:=\begin{bmatrix}q_{2}&x\\ x^{*}&q_{3}\end{bmatrix},\quad B_{3}:=\begin{bmatrix}pq_{2}&s_{23}\\ s_{23}^{*}&pq_{3}\end{bmatrix}\geq 0,$ (43)
$B_{4}:=\begin{bmatrix}(1-p)q_{2}&-s_{23}+x\\ (-s_{23}+x)^{*}&(1-p)q_{3}\end{bmatrix}\geq 0.$ (44)
Note that $B_{3}+B_{4}=D_{1}$ for $B_{3}$$B_{4}$ in (43). From $|D_{1}|=0$, there is a nonzero vector $v=[a,b]^{\dagger}$ such that $D_{1}v=0$. That is,
$q_{2}a+xb=0,\quad x^{*}a+q_{3}b=0.$ (45)
Then we have $B_{3}v+B_{4}v=0$, and hence $v^{*}B_{1}v+v^{*}B_{2}v=0$. By $B_{3},B_{4}\geq 0$, we have $v^{*}B_{3}v\geq 0$ and $v^{*}B_{4}v\geq 0$. Then $v^{*}B_{3}v=v^{*}B_{4}v=0$, and thus $B_{3}v=B_{4}v=0$. Using (45), we have
$B_{3}v=-B_{4}v=[b(s_{23}-px),a(s_{23}^{*}-px^{*})]^{\dagger}=0$ (46)
From the first equality of (45), we have $a\neq 0$ or $b\neq 0$ in $v$. By (46), we have $s_{23}=px$. Next we show $s_{27}=s_{36}=0$. Since the following principal minors are positive definite, i.e.
$\begin{bmatrix}pq_{2}&s_{23}&s_{27}\\ s_{23}^{*}&pq_{3}&0\\ s_{27}^{*}&0&(1-p)q_{3}\end{bmatrix},\begin{bmatrix}pq_{2}&s_{23}&0\\ s_{23}^{*}&pq_{3}&s_{36}\\ 0&s_{36}^{*}&(1-p)q_{2}\end{bmatrix}$ (47)
we have $(1-p)q_{3}|D_{1}|-|s_{27}|^{2}pq_{3}\geq 0$, $(1-p)q_{2}|D_{1}|-|s_{36}|^{2}pq_{2}\geq 0$. By $|D_{1}|=0$, we obtain that $s_{27}=s_{36}=0$. To conclude, we show that $s_{23}=px$ and $s_{27}=s_{36}=0$, i.e. $\sigma=\rho$, if $|x|=\sqrt{q_{2}q_{3}}$. We have proven the ”if” part.
We show the ”only if” part. Suppose $|D_{1}|>0$. We show that there is another state $\sigma_{1}$ compatible with $\rho$ by all possible 2-marginals. In detail,we set all elements of $\sigma_{1}$ the same as $\rho$, except for $s_{27}$. Now we show that the element $s_{27}$ of $\sigma_{1}$ is not necessarily equal to zero. We check all principal minors of $\sigma_{1}$ that contains the element $s_{27}$, and obtain that
$0<|s_{27}|^{2}\leq p(1-p)(q_{2}q_{2}-|x|^{2})^{2}/q_{2}q_{3}.$ (48)
Note that the RHS of (48) is strictly larger than zero. Hence we can choose $s_{27}>0$ satisfying (48) in $\sigma_{1}$. One can verify that $\sigma_{1}$ is compatible with $\rho$ by all possible 2-marginals and $\sigma_{1}\neq\rho$. Hence $\rho$ is not 2-UDA. This completes the proof.   $\sqcap\sqcup$
The proof of Proposition 13 is presented, where we show that a state $\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}$ with $\mathop{\rm rank}(\rho_{A_{3}A_{4}})$ is rank-three and four is not 2-UDA.
Proof.
(i) First we claim that
$\ker(\rho_{A_{2}A_{3}})=\text{span}\{\cos\theta|00\rangle+\sin\theta|11\rangle,\;\theta\in(0,\frac{\pi}{4}]\}.$
In fact, we have $\mathop{\rm Dim}\ker(\rho_{A_{2}A_{3}})=1$. Suppose a state $|\beta\rangle\in\ker(\rho_{A_{2}A_{3}})$. If $|\beta\rangle$ is a product pure state, then $|\beta\rangle$ is LU equivalent to $|00\rangle$, which implies that $\rho_{A_{2}A_{3}}$ has a zero diagonal element. This shows the claim. Hence, the range of $\rho_{A_{2}A_{3}}$ is given by
${\cal R}(\rho_{A_{2}A_{3}})=\text{span}\{\sin\theta|00\rangle-\cos\theta|11\rangle,\;|jk\rangle:jk\neq 00,11\}.$
Suppose $\rho_{A_{1}}=\mathop{\rm diag}\{p_{1},p_{2},...,p_{d_{1}}\}$. Let
$\sigma_{A_{1}A_{2}A_{3}}=\begin{bmatrix}p_{1}\rho_{A_{2}A_{3}}&0&\cdots&0&B\\ 0&p_{2}\rho_{A_{2}A_{3}}&\cdots&0&0\\ \vdots&\vdots&&\vdots&\vdots\\ B^{\dagger}&0&\cdots&0&p_{d_{1}}\rho_{A_{2}A_{3}}\end{bmatrix}$, (49)
where $B$ is a matrix where only the bottom-left element is $x$, and $x>0$ small enough. One can verify that the state $\sigma$ is compatible with $\rho$ by all 2-marginals. We show that $\sigma\geq 0$. From ${\cal R}(B)\subset{\cal R}(\rho_{A_{2}A_{3}})$, the state $\sigma$ is congruent to the matrix
$(p_{1}\rho_{A_{2}A_{3}}-BV^{\dagger}VB^{\dagger})\oplus_{k=2}^{d_{1}-1}p_{2}\rho_{A_{2}A_{3}}\oplus I_{d_{2}d_{3}-1}\oplus 0.$
It suffices to show that
$\xi:=$ $p\rho_{A_{2}A_{3}}-BV^{\dagger}VB^{\dagger}$ (50)
$=$ $p\rho_{A_{2}A_{3}}-\mathop{\rm diag}\{0,0,\cdots,0,y\}\geq 0,$
where $y\rightarrow 0$ when $x\rightarrow 0$. By Lemma 3, there is $x>0$ small enough such that $\xi\geq 0$, and hence $\sigma\geq 0$. This completes the proof.
(ii) If $\mathop{\rm rank}(\rho_{A_{2}A_{3}})=4$, then the state $\sigma$ in (49) is also compatible with $\rho$. Further $\sigma\geq 0$ holds since (50) is derived by ${\cal R}(\mathop{\rm diag}\{0,0,\cdots,0,y\})\subseteq{\cal R}(\rho_{A_{2}A_{3}})$ and Lemma 3.   $\sqcap\sqcup$
Appendix C Proof four-qubit UDA states
Firstly, Proposition 18 shows the characterization of the tensor product of rank-two two-qubit states. The proof of Proposition 18 is given here.
Proof.
(i) By some observations, we obtain that $\rho_{A_{1}A_{2}}=\rho_{A_{1}}\otimes\rho_{A_{2}}$ iff $p_{2}p_{3}=0$. In this case, $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}}\otimes\rho_{A_{3}A_{4}}$ has been investigated in Proposition 17. The same analysis applies to $\rho_{A_{2}A_{3}}$.
(ii) We show the proof by two cases: (ii.a) $q_{1}=0$ and (ii.b) $q_{1}>0$. (ii.a) If $p_{1}=0$, $|a|<\sqrt{p_{2}p_{3}}$ and $q_{1}=0$, $|x|<\sqrt{q_{2}q_{3}}$, then $\rho$ is not 2-UDA. We establish a state $\sigma$ compatible with $\rho$, where
$\sigma=0_{5\times 5}\oplus H\oplus 0_{5\times 5}.$ (51)
For $|\epsilon|>0$ small enough, $H=\begin{bmatrix}H_{11}&0&H_{12}\\ 0&0&0\\ H_{12}^{\dagger}&0&H_{22}\end{bmatrix}$ for $H_{11}=\begin{bmatrix}p_{2}q_{2}&p_{2}x\\ (p_{2}x)^{*}&p_{2}q_{3}\end{bmatrix}$, $H_{12}=\begin{bmatrix}aq_{2}&ax+\epsilon\\ (ax)^{*}&aq_{3}\end{bmatrix}$ and $H_{22}=\begin{bmatrix}p_{3}q_{2}&p_{3}x\\ (p_{3}x)^{*}&p_{3}q_{3}\end{bmatrix}$. One can verify that $H\geq 0$. In fact, by discarding the zero rows and columns of $H$, the remaining matrix is given by $A=\begin{bmatrix}p_{2}&a\\ a^{*}&q_{2}\end{bmatrix}\otimes\begin{bmatrix}q_{2}&x\\ x^{*}&q_{3}\end{bmatrix}>0$. Then the perturbation remains the positive semi-definiteness of $A$. Since $\epsilon\neq 0$, we have $\sigma\neq\rho$. Therefore, $\rho$ is not 2-UDA.
(ii.b) $p_{1}=0$, $|a|<\sqrt{p_{2}p_{3}}$ and $q_{1}q_{2}q_{3}>0$ (equivalently, $p_{1}p_{2}p_{3}>0$ and $q_{1}=0$, $|x|<\sqrt{q_{2}q_{3}}$). In this case, the state $\rho$ is given by
$\rho=$ $\rho_{A_{1}A_{2}}\otimes\rho_{A_{3}A_{4}}$
$=$ $\begin{bmatrix}0&0&0&0\\ 0&p_{2}&a&0\\ 0&a^{*}&p_{3}&0\\ 0&0&0&0\end{bmatrix}\otimes\begin{bmatrix}q_{1}&y&z&0\\ y^{*}&q_{2}&x&0\\ z^{*}&x^{*}&q_{3}&0\\ 0&0&0&0\end{bmatrix}$.
Since $|a|<\sqrt{p_{2}p_{3}}$, there is a $D=\mathop{\rm diag}\{0,\epsilon_{1},\epsilon_{2},0\}$ with $\epsilon_{k}>0$ small enough, such that
$\rho=(\rho_{A_{1}A_{2}}-D)\otimes\rho_{A_{3}A_{4}}+\alpha,$ (52)
where $\alpha$ in (9) is not 2-UDA, and $\rho_{A_{1}A_{2}}-D\geq 0$. Using Lemma 5 and (52), $\rho$ is not 2-UDA.
(iii) We show that $\rho$ is 2-UDA by contradiction. Suppose $\rho$ is not 2-UDA and $\sigma$ is a state compatible with $\rho$. Let $\chi:=\rho-\sigma$. Then $\chi$ is a Hermitian operator $\chi$ on ${\cal H}_{A_{1}A_{2}A_{3}A_{4}}$ satisfying $\mathop{\rm Tr}_{A_{j}A_{k}}\chi=0$. Obviously, $\rho$ is not 2-UDA iff $\chi\neq 0$ for some state $\sigma$. However, one can show that $\chi=0$ as follows. From Lemma 6, we have ${\cal R}(\sigma)\subseteq{\cal R}(\rho)$, and hence
${\cal R}(\chi)\subseteq{\cal R}(\rho)=\text{span}\{|a_{i}\rangle:i=1,2,3,4\}.$ (53)
Let $A=(|a_{1}\rangle,|a_{2}\rangle,|a_{3}\rangle,|a_{4}\rangle)$ be a matrix with rank equal to four. Since $\chi$ is Hermitian and ${\cal R}(\chi)\subseteq{\cal R}(\rho)$, we claim that without loss of generality,
$\chi=ASA^{\dagger},$ (54)
where $S=(s_{kl})\in\mathbb{C}^{4\times 4}$ is a Hermitian matrix for $k,l=1,2,3,4$. From $\chi=\chi^{\dagger}$, we have $S=:(A^{\dagger}A)^{-1}A^{\dagger}\chi A(A^{\dagger}A)^{-1}$ is Hermitian. One can verify that $ASA^{\dagger}=\chi$ as follows. First $P_{A}=A(A^{\dagger}A)^{-1}A^{\dagger}$ is a projector on ${\cal R}(A)={\cal R}(\rho)$ since $P_{A}x=x$, $\forall x\in{\cal R}(A)$ and $P_{A}$ is self-adjoint and idempotent. From ${\cal R}(\chi)\subset{\cal R}(\rho)$, we have $P_{A}\chi=\chi$, and $\chi P_{A}=\chi$ by Hermitian $\chi$. This implies that $ASA^{\dagger}=P_{A}\chi P_{A}=\chi$. This proves the claim. Note that $p_{1}p_{2}p_{3}>0$ and $q_{1}q_{2}q_{3}>0$. Up to LU equivalence, the range of $\rho_{A_{i}A_{i+1}}$, for $i=1,3$, is one of the following
${\cal R}(\rho_{A_{i}A_{i+1}})=\text{span}\{m_{i}|00\rangle+n_{i}|01\rangle,\;g_{i}|00\rangle+h_{i}|10\rangle\},$
${\cal R}(\rho_{A_{i}A_{i+1}})=\text{span}\{z_{i}|00\rangle,\;u_{i}|01\rangle+v_{i}|10\rangle\},$ (56)
where $n_{i}h_{i}\neq 0$ and $z_{i}u_{i}v_{i}\neq 0$ since $\rho_{A_{i}A_{i+1}}$ has two diagonal elements if they are zero, and $z_{i}\neq 0$ by $\mathop{\rm rank}(\rho_{A_{i}A_{i+1}})=2$.
The proof of $\chi=0$ above consists of three parts (iii.a)-(iii.c), depending on the range of $\rho_{A_{1}A_{2}}$ and $\rho_{A_{3}A_{4}}$.
(iii.a) ${\cal R}(\rho_{A_{1}A_{2}})$ and ${\cal R}(\rho_{A_{3}A_{4}})$ follow (C). Then $|a_{i}\rangle$’s in (53) are
$\hskip-14.22636pt|a_{1}\rangle=(m_{1}|00\rangle+n_{1}|01\rangle)\otimes(m_{3}|00\rangle+n_{3}|01\rangle),$ (57)
$\hskip-14.22636pt|a_{2}\rangle=(m_{1}|00\rangle+n_{1}|01\rangle)\otimes(g_{3}|00\rangle+h_{3}|10\rangle),$ (58)
$\hskip-14.22636pt|a_{3}\rangle=(g_{1}|00\rangle+h_{1}|10\rangle)\otimes(m_{3}|00\rangle+n_{3}|01\rangle),$ (59)
$\hskip-14.22636pt|a_{4}\rangle=(g_{1}|00\rangle+h_{1}|10\rangle)\otimes(g_{3}|00\rangle+h_{3}|10\rangle).$ (60)
From $\mathop{\rm rank}(\rho_{A_{1}A_{2}})=\mathop{\rm rank}(\rho_{A_{3}A_{4}})=2$, at most one element of the two sets $\{m_{i},g_{i}\}$ is zero, respectively.
We show that $\mathop{\rm Tr}_{A_{j}A_{k}}\chi=0$ implies $S=0$ as follows. Let $\chi=(H_{ij})$, where $H_{i,j}\in\mathbb{C}^{2\times 2}$ and i,j=1,2,,8i,j=1,2,...,8. By (54), we have $H_{j,k}=0$ if $j,k=7,8$. From $\chi_{A_{1}A_{2}}=\chi_{A_{1}A_{3}}=\chi_{A_{2}A_{3}}=0$ and $H_{77}=H_{88}=0$, we have $\mathop{\rm Tr}{H_{jj}}=0$ for j=1,2,,8j=1,2,...,8. In particular, we have $s_{22}=s_{44}=0$ by $\mathop{\rm Tr}(H_{44})=|h_{3}|^{2}|n_{1}|^{2}s_{22}=0$ and $\mathop{\rm Tr}(H_{66})=|h_{1}|^{2}|h_{3}|^{2}s_{44}=0$. Next we show that $s_{34}=0$. From $s_{22}=s_{44}=0$ and $\chi_{A_{1}A_{4}}=0$, we have $H_{55}=0$, and thus $|h_{1}|^{2}g_{3}^{*}n_{3}s_{34}=0$. By $\chi_{A_{1}A_{3}}=0$, we have $\mathop{\rm Tr}(H_{56})=|h_{1}|^{2}h_{3}^{*}m_{3}s_{34}=0$. Since $m_{3}$ and $g_{3}$ cannot be zero simultaneously, we have $s_{34}=0$. Next, we have $s_{24}=0$ since $\chi_{A_{1}A_{3}}=0$ and $\chi_{A_{2}A_{3}}=0$ derive that $\mathop{\rm Tr}(H_{26})=h_{1}^{*}|h_{3}|^{2}m_{1}s_{24}=0$ and $\mathop{\rm Tr}(H_{24})=|h_{3}|^{2}n_{1}^{*}g_{1}s_{24}=0$. Note that $m_{1}$ and $g_{1}$ cannot be zero at the same time. Hence $s_{24}=0$. Besides, $\chi_{A_{1}A_{4}}=\chi_{A_{2}A_{4}}=0$ derives that $H_{33}=H_{55}=0$, and thus $|n_{1}|^{2}|n_{3}|^{2}s_{11}=|h_{1}|^{2}|n_{3}|^{2}s_{33}=0$. We have $s_{11}=s_{33}=0$. Using $\chi_{A_{1}A_{4}}=\chi_{A_{2}A_{4}}=0$, we have $H_{15}=H_{13}=0$, and thus $|n_{3}|^{2}h_{1}^{*}m_{1}s_{13}=|n_{3}|^{2}n_{1}g_{1}^{*}s_{13}=0$. We obtain that $s_{13}=0$. Next we show that $s_{12}=0$. By $\chi_{A_{2}A_{4}}=\chi_{A_{2}A_{3}}=0$, we have $H_{33}=0$ and $\mathop{\rm Tr}(H_{34})=0$, which implies $|n_{1}|^{2}n_{3}g_{3}^{*}s_{12}=|n_{1}|^{2}h_{3}^{*}m_{3}s_{12}=0$. As a complicated case, we show that $s_{14}=0$. From $\chi_{A_{2}A_{3}}=0$, we have $\mathop{\rm Tr}(H_{14})=h_{3}^{*}n_{1}^{*}g_{1}m_{3}s_{14}=0$, $\mathop{\rm Tr}(H_{23})=h_{3}^{*}n_{1}g_{1}^{*}m_{3}s_{14}=0$ and $H_{13}=0$, which derives $n_{1}n_{3}g_{1}^{*}g_{3}^{*}s_{14}=0$. Using $\chi_{A_{1}A_{4}}=0$, we have $H_{15}=0$ and thus $n_{3}h_{1}^{*}m_{1}g_{3}^{*}s_{14}=0$. By $\chi_{A_{1}A_{3}}=0$, we have $\mathop{\rm Tr}{(H_{16})}=h_{1}^{*}h_{3}^{*}m_{1}m_{3}s_{14}=0$. Note that at most one element of the two sets $\{m_{i},g_{i}\}$ is zero, respectively. Hence $s_{14}=0$. Finally, we show that $s_{23}=0$. From $\chi_{A_{2}A_{4}}=\chi_{A_{1}A_{4}}=0$, we have $H_{13}=H_{15}=0$, and thus $n_{1}n_{3}^{*}g_{1}^{*}g_{3}s_{23}=h_{1}^{*}n_{3}^{*}m_{1}g_{3}s_{23}=0$. By $\chi_{A_{1}A_{3}}=\chi_{A_{2}A_{3}}=0$, we have $\mathop{\rm Tr}(H_{25})=h_{1}^{*}h_{3}m_{1}m_{3}^{*}s_{23}=0$ and $\mathop{\rm Tr}(H_{14})=h_{3}n_{1}g_{1}^{*}m_{3}^{*}s_{23}=0$. We have $s_{23}=0$. To conclude, we show that the Hermitian operator $S=0$, and thus $\chi=\rho-\sigma=0$ for any $\sigma$. It is a contraction. Therefore, $\rho$ is 2-UDA.
(iii.b) ${\cal R}(\rho_{A_{1}A_{2}})$ and ${\cal R}(\rho_{A_{3}A_{4}})$ follow (56). By the same analysis as (iii.a), one can verify that $S=0$ and thus $\chi=0$ for any $\sigma$, which is shown in TABLE 2. The state $\rho$ is 2-UDA in this case.
(iii.c) ${\cal R}(\rho_{A_{1}A_{2}})$ follows (56) and ${\cal R}(\rho_{A_{3}A_{4}})$ follows (C), up to system permutation of $A_{1}A_{2}$ and $A_{3}A_{4}$. We obtain that $S=0$ in TABLE 2, and $\chi=ASA^{\dagger}=0$. Therefore, $\rho$ is 2-UDA. This completes the proof.   $\sqcap\sqcup$
${\cal R}(\rho_{A_{1}A_{2}})$ and ${\cal R}(\rho_{A_{3}A_{4}})$ follow (56). ${\cal R}(\rho_{A_{1}A_{2}})$ follows (56) and ${\cal R}(\rho_{A_{3}A_{4}})$ follows (C)
$\mathop{\rm Tr}(H_{22})=|v_{3}|^{2}|z_{1}|^{2}s_{22}=0\Rightarrow s_{22}=0$ $\mathop{\rm Tr}(H_{22})=|h_{3}|^{2}|z_{1}|^{2}s_{22}=0\Rightarrow s_{22}=0$
$\mathop{\rm Tr}(H_{44})=|v_{3}|^{2}|u_{1}|^{2}s_{44}=0\Rightarrow s_{44}=0$ $\mathop{\rm Tr}(H_{66})=|v_{1}|^{2}|h_{1}|^{2}s_{44}=0\Rightarrow s_{44}=0$
$\mathop{\rm Tr}(H_{14})=u_{1}^{*}v_{3}^{*}z_{1}z_{3}s_{14}=0\Rightarrow s_{14}=0$ $\mathop{\rm Tr}(H_{26})=|h_{3}|^{2}v_{1}^{*}z_{1}s_{24}=0\Rightarrow s_{24}=0$
$\mathop{\rm Tr}(H_{26})=|v_{3}|^{2}v_{1}^{*}z_{1}s_{24}=0\Rightarrow s_{24}=0$ $H_{15}=0\Rightarrow|n_{3}|^{2}z_{1}v_{1}^{*}s_{13}=0\Rightarrow s_{13}=0$
$\mathop{\rm Tr}(H_{25})=v_{1}^{*}v_{3}z_{1}z_{3}^{*}s_{23}=0\Rightarrow s_{23}=0$ $\left.\begin{array}[]{l}\mathop{\rm Tr}(H_{14})=u_{1}^{*}h_{3}^{*}z_{1}m_{3}s_{14}=0\\ H_{13}=0\Rightarrow u_{1}^{*}z_{1}n_{3}g_{3}^{*}s_{14}=0\end{array}\right\}\Rightarrow s_{14}=0$
$\mathop{\rm Tr}(H_{35})=u_{1}v_{1}^{*}|z_{3}|^{2}s_{33}=0\Rightarrow s_{33}=0$ $\left.\begin{array}[]{l}\mathop{\rm Tr}(H_{25})=v_{1}^{*}h_{3}z_{1}m_{3}^{*}s_{23}=0\\ H_{15}=0\Rightarrow n_{3}^{*}v_{1}^{*}z_{1}g_{3}s_{23}=0\end{array}\right\}\Rightarrow s_{23}=0$
$H_{11}=0\Rightarrow|z_{1}|^{2}|z_{3}|^{2}s_{11}=0\Rightarrow s_{11}=0$ $H_{55}=0\Rightarrow|n_{3}|^{2}|v_{1}|^{2}s_{33}=0\Rightarrow s_{33}=0$
$H_{11}=0\Rightarrow|z_{1}|^{2}z_{3}u_{3}^{*}s_{12}=0\Rightarrow s_{12}=0$ $H_{11}=0\Rightarrow|n_{3}|^{2}|z_{1}|^{2}s_{11}=0\Rightarrow s_{11}=0$
$H_{13}=0\Rightarrow|z_{3}|^{2}z_{1}u_{1}^{*}s_{13}=0\Rightarrow s_{13}=0$ $\left.\begin{array}[]{l}\mathop{\rm Tr}(H_{34})=|u_{1}|^{2}n_{3}^{*}g_{3}s_{34}=0\\ H_{33}=0\Rightarrow|u_{1}|^{2}h_{3}^{*}m_{3}s_{34}=0\end{array}\right\}\Rightarrow s_{34}=0$
$H_{55}=0\Rightarrow|v_{1}|^{2}z_{3}u_{3}^{*}s_{34}=0\Rightarrow s_{34}=0$ $\left.\begin{array}[]{l}H_{11}=0\Rightarrow|z_{1}|^{2}n_{3}g_{3}^{*}s_{12}=0\\ \mathop{\rm Tr}(H_{12})=|z_{1}|^{2}h_{3}^{*}m_{3}s_{12}=0\end{array}\right\}\Rightarrow s_{12}=0$
Table 2: The elements $s_{kl}$ of $S$ given in (54), under the condition $\mathop{\rm Tr}_{A_{j}A_{k}}\chi=0$, where $\chi=(H_{ij})$ and $H_{i,j}\in\mathbb{C}^{2\times 2}$ for i,j=1,2,,8i,j=1,2,...,8
Proposition 19 shows that the tensor product of two mixed states, one of which is not pure and the rank of the other one is larger than two, is not 2-UDA. The proof is shown as follows.
Proof.
As the first step, we assume that $\mathop{\rm rank}(\rho_{A_{1}A_{2}})=2$ and $\mathop{\rm rank}(\rho_{A_{3}A_{4}})=3$. Then $\rho$ is given by
$\rho=\rho_{A_{1}A_{2}}\otimes\rho_{A_{3}A_{4}}:=\begin{bmatrix}p_{1}&b&c&0\\ b^{*}&p_{2}&a&0\\ c^{*}&a^{*}&p_{3}&0\\ 0&0&0&0\end{bmatrix}\otimes\rho_{A_{3}A_{4}}.$ (61)
The UDA property of $\rho$ can be analyzed by the following cases.
(i) If $p_{2}p_{3}=0$, then $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}}\otimes\rho_{A_{3}A_{4}}$ and $\rho$ is not 2-UDA since $\rho_{A_{2}}\otimes\rho_{A_{3}A_{4}}$ is not 2-UDA.
(ii) $p_{1}=0$ and $p_{2}p_{3}>0$; Let
$D=\mathop{\rm diag}\{0,\epsilon_{1},\epsilon_{2},0\}$ (62)
with $\epsilon_{i}>0$ small enough. We show that $\rho$ is not 2-UDA since $\rho_{A_{1}A_{2}}=(\rho_{A_{1}A_{2}}-D)+D:=\alpha_{1}+\alpha_{2}$, for $\alpha_{1}\geq 0$. By similar analysis as (9), we obtain that $\alpha_{2}\otimes\rho_{A_{2}A_{4}}$ is not 2-UDA. In detail, a state $\sigma$ compatible with $\rho$ is given by $\sigma=0_{2}\oplus\begin{bmatrix}\epsilon_{1}\rho_{A_{3}A_{4}}&B\\ B^{\dagger}&\epsilon_{2}\rho_{A_{3}A_{4}}\end{bmatrix}\oplus 0_{2}$ with $\mathop{\rm Tr}B=0$. We can choose $B=\mathop{\rm diag}\{0,x,-x,0\}$ for $x>0$ small enough, such that ${\cal R}(B)\subset{\cal R}(\rho_{A_{3}A_{4}})$. Then $\sigma\neq\rho$ is compatible with $\rho$. From Lemma 5, $\rho$ is not 2-UDA.
(iii) $p_{1}p_{2}p_{3}>0$. Since $\rho_{A_{2}A_{4}}$ is rank-three, we claim that up to LU equivalence, ${\cal R}(D)\subseteq{\cal R}(\rho_{A_{3}A_{4}})$, where $D$ is given in (62). In fact, if $\rho_{A_{3}A_{4}}$ is LU equivalent to a state with a zero diagonal element, we have $\rho_{A_{3}A_{4}}=N\oplus 0$ without loss of generality, where $N$ is full rank. The claim holds. Otherwise, $\rho_{A_{3}A_{4}}$ is not LU equivalent to the state with a zero diagonal element. The range of $\rho_{A_{2}A_{3}}$ is ${\cal R}(\rho_{A_{3}A_{4}})=\text{span}\{|\psi_{1}\rangle:=|01\rangle,|\psi_{2}\rangle:=|10\rangle,|\psi\rangle:=\sin\theta|00\rangle-\cos\theta|11\rangle\}$. We have proven the claim. Then we have $\rho=\rho_{A_{1}A_{2}}\otimes(\rho_{A_{3}A_{4}}-D)+\rho_{A_{1}A_{2}}\otimes D:=\beta_{1}+\beta_{2}$. We showed that $\beta$ is not 2-UDA in the analysis below (9). Using Lemme 5, $\rho$ is not 2-UDA.
Next we show that $\rho$ is not 2-UDA if $\mathop{\rm rank}(\rho_{A_{1}A_{2}})=$ and $\mathop{\rm rank}(\rho_{A_{3}A_{4}})=4$. The spectral decomposition of $\rho_{A_{3}A_{4}}=(\sum_{i=1}^{3}\lambda_{i}|a_{j}\rangle\!\langle a_{j}|)+\lambda_{4}|a_{4}\rangle\!\langle a_{4}|:=p\alpha+(1-p)\beta$, where $p=\sum_{i=1}^{3}\lambda_{i}$ and $\alpha=\frac{1}{p}\sum_{i=1}^{3}\lambda_{i}|a_{j}\rangle\!\langle a_{j}|$ is rank-three. Then $\rho=p\rho_{A_{1}A_{2}}\otimes\alpha+(1-p)\rho_{A_{1}A_{2}}\otimes\beta$, where the first term $\rho_{A_{1}A_{2}}\otimes\alpha$ is not 2-UDA by the preceding analysis. Using Lemma 5, $\rho$ is not 2-UDA. Similarly, one can show that $\rho$ is not 2-UDA if $\mathop{\rm rank}(\rho_{A_{1}A_{2}})\geq 3$ and $\mathop{\rm rank}(\rho_{A_{3}A_{4}})\geq 3$. The UDA property is invariant up to system permutations between systems $A_{1}A_{2}$ and $A_{3}A_{4}$. This completes the proof.   $\sqcap\sqcup$
Appendix D Proof of $n$-qubit UDA states
Lemma 21 shows that $\rho=\rho_{A_{1}}\otimes(D\oplus 0)\otimes(H\oplus 0)$ with mixed $\rho_{A_{1}}$ is not 2-UDA. The proof is presented here.
Proof.
Suppose $\sigma$ is a state 2-compatible with $\rho$. By Lemma 5 (iii), $\sigma$ further satisfies
$\sigma_{A_{2}A_{3}A_{4}A_{5}}=\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}},$ (65)
$\sigma_{A_{1}A_{2}A_{3}}=\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}},$
$\sigma_{A_{1}A_{4}A_{5}}=\rho_{A_{1}}\otimes\rho_{A_{4}A_{5}}.$
Obviously, these conditions imply that $\sigma$ is 2-compatible with $\rho$. From (65), any $\sigma$ can be given by
$\sigma_{A_{1}...A_{5}}=\begin{bmatrix}X_{A_{2}...A_{5}}&Y_{A_{2}...A_{5}}\\ Y_{A_{2}...A_{5}}^{\dagger}&\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}}-X_{A_{2}...A_{5}}\end{bmatrix}$, (66)
where by (65) and (65), we have $0\leq X\leq\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}}$ $X_{A_{2}A_{3}}=h_{1}\rho_{A_{2}A_{3}},\;X_{A_{4}A_{5}}=h_{1}\rho_{A_{4}A_{5}}$, and $Y_{A_{2}A_{3}}=Y_{A_{4}A_{5}}=0$. Then $\rho$ is 2-UDA iff $\sigma=\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}}$ iff $X=h_{1}\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}}$ and $Y=0$. We establish a different $\sigma$ that is compatible with $\rho$, such that $X=h_{1}\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}}$ but $Y\neq 0$. Then
$\sigma=\begin{bmatrix}h_{1}\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}}&Y\\ Y^{\dagger}&h_{2}\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}}\end{bmatrix}$. (67)
In order to make $\sigma\geq 0$, we choose nonzero Hermitian $Y$ satisfying ${\cal R}(Y)\subseteq{\cal R}(\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}})$, $Y_{A_{2}A_{3}}=Y_{A_{4}A_{5}}=0$ and $\lVert Y\rVert$ small enough. By the same analysis as Lemma 16, $\sigma$ is congruent to $I_{4}\oplus 0_{12}\oplus(h_{2}\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}}-Y^{\dagger}Y)$, which is positive semidefinite. We show the expression of $Y$. Since ${\cal R}(Y)\subseteq{\cal R}(\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}})$, we have $Y=ASA^{\dagger}$, where $S=(s_{jk})$ for $j,k=1,2,3,4$ is Hermitian, and $A=(|a_{1}\rangle,|a_{2}\rangle,|a_{3}\rangle,|a_{4}\rangle)$. Here $|a_{i}\rangle$’s are the base of the vector space ${\cal R}(\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}})$. Depending on the range of ${\cal R}(\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}})$, this is derived by three cases. First, ${\cal R}(\rho_{A_{2}A_{3}})$ and ${\cal R}(\rho_{A_{4}A_{5}})$ both follow (C). Then $|a_{i}\rangle$’s are given in (57)-(60), and $S$ is diagonal with $s_{11}=x$, $s_{22}=-\frac{|m_{3}|^{2}+|n_{3}|^{2}}{|g_{3}|^{2}+|h_{3}|^{2}}x$, $s_{33}=-\frac{|m_{1}|^{2}+|n_{1}|^{2}}{|g_{1}|^{2}+|h_{1}|^{2}}x$, $s_{44}=-\frac{(|m_{1}|^{2}+|n_{1}|^{2})(|m_{3}|^{2}+|n_{3}|^{2})}{(|g_{1}|^{2}+|h_{1}^{2}|)(|g_{3}|^{2}+|h_{3}^{2}|)}x$, for $x>0$ small enough. Second, ${\cal R}(\rho_{A_{2}A_{3}})$ and ${\cal R}(\rho_{A_{4}A_{5}})$ both follow (56). Then $|a_{1}\rangle=z_{1}z_{3}|00\rangle$, $|a_{2}\rangle=z_{1}|00\rangle\otimes(u_{3}|01\rangle+v_{3}|10\rangle)$, $|a_{3}\rangle=(u_{1}|01\rangle+v_{1}|10\rangle)\otimes z_{3}|00\rangle$, $|a_{4}\rangle=(u_{1}|01\rangle+v_{1}|10\rangle)\otimes(u_{3}|01\rangle+v_{3}|10\rangle)$. We choose $s_{12}=(|u_{1}|^{2}+|v_{1}|^{2})x$, $s_{34}=-|z_{1}|^{2}x$, and other elements of $S$ are zero, for $x>0$ small enough. Third, up to system permutation of $A_{2}A_{3}$ and $A_{4}A_{5}$, ${\cal R}(\rho_{A_{2}A_{3}})$ and ${\cal R}(\rho_{A_{4}A_{5}})$ follow (56) and (C), respectively. Then $|a_{1}\rangle=z_{1}|00\rangle\otimes(m_{3}|00\rangle+n_{3}|01\rangle)$, $|a_{2}\rangle=z_{1}|00\rangle\otimes(g_{3}|00\rangle+h_{3}|10\rangle)$, $|a_{3}\rangle=(u_{1}|01\rangle+v_{1}|10\rangle)\otimes(m_{3}|00\rangle+n_{3}|01\rangle)$, and $|a_{4}\rangle=(u_{1}|01\rangle+v_{1}|10\rangle)\otimes(g_{3}|00\rangle+h_{3}|10\rangle)$. We choose $s_{13}=-(|g_{3}|^{2}+|h_{3}|^{2})x$, $s_{24}=(|m_{3}|^{2}+|n_{3}|^{2})x$ and other elements of $S$ are zero, for $x>0$ small enough. To conclude, we can always find a nonzero $Y$ in (67) such that $\sigma$ is 2-compatible with $\rho$. This completes the proof.   $\sqcap\sqcup$
Proposition 23 characterizes some $n$-qubit $k$-UDA states. Its proof is given as follows.
Proof.
(i) The ”if part” is trivial by Lemma 2. We show the ”only if” part as follows. Suppose there are only nk1n-k-1 of the $\rho_{A_{i}}$’s that are pure. From Lemma 7, without loss of generality, we assume that $\rho_{A_{j}}=|\psi^{(j)}\rangle\!\langle\psi^{(j)}|$ for j=1,2,,nk1j=1,2,...,n-k-1 are pure. For any state $\sigma$ $k$-compatible with $\rho$, we have $\sigma=\otimes_{i=1}^{n-k-1}|\psi^{(j)}\rangle\!\langle\psi^{(j)}|_{A_{j}}\otimes\sigma_{A_{n-k},...,A_{n}}$. It suffices to determine whether the $(k+1)$-partite state $\sigma_{A_{n-k},...,A_{n}}$ is $\otimes_{i=n-k}^{n}\rho_{A_{i}}$ for any $\sigma$ and mixed state $\rho_{A_{i}}$. We show that a $(k+1)$-partite state $\lambda=\otimes_{i=1}^{k+1}\rho_{A_{i}}$ is not $k$-UDA by induction. The case for $k=1$ holds by Proposition 8. Suppose the fact holds for $k$-partite state. We show that it is also hold for a $k+1$-partite state $\lambda=\otimes_{i=1}^{k+1}\rho_{A_{i}}$. Since $\lambda_{A_{1}...A_{k}}=\otimes_{i=1}^{k}\rho_{A_{i}}$ is not (k1)(k-1)-UDA, there is a different state $\sigma^{\prime}$ that is (k1)(k-1)-compatible with $\lambda_{A_{1}...A_{k}}$. Then $\sigma=\sigma^{\prime}\otimes\rho_{A_{k+1}}$ is $k$-compatible with $\lambda$, and $\sigma\neq\lambda$. This proves the claim. Hence, $\sigma_{A_{n-k},...,A_{n}}$ is not necessarily $\otimes_{i=n-k}^{n}\rho_{A_{i}}$.
(ii) From Lemma 4, $\rho$ is 2-UDA only if $\rho_{A_{n-1}}\otimes\rho_{A_{n-1}A_{n}}$ is 2-UDA. Using Theorem 14, this derives four cases, (a) $\rho_{A_{n-1}A_{n}}$ is 2-UDA. Then $\rho$ is 2-UDA iff $\otimes_{i=1}^{n-2}\rho_{A_{i}}$ is 2-UDA; (b) $\rho_{A_{n-1}}$ or $\rho_{A_{n}}$ is pure. By Lemma 2, $\rho=\otimes_{i=1}^{n}\rho_{A_{i}}$, which is considered in (i); (c) $\rho_{A_{n-2}}$ is pure. If $\otimes_{i=1}^{n-2}\rho_{A_{i}}$ is pure, then $\rho$ is 2-UDA for any $\rho_{A_{n-1}}\rho_{A_{n}}$; (d) if less than n2n-2 states of $\rho_{A_{i}}$ are pure for i=1,2,,n2i=1,2,...,n-2, then $\rho$ is 2-UDA only when $\rho_{A_{n-1}A_{n}}$ is rank-two and LU equivalent to $D\oplus 0$. Otherwise, there is a mixed $\rho_{A_{k}}$ for k=1,2,,n2k=1,2,...,n-2 such that $\rho_{A_{k}}\otimes\rho_{A_{n-1}A_{n}}$ is not 2-UDA. We claim that $\rho=\otimes_{i=1}^{n-2}\rho_{A_{i}}\otimes(D\oplus 0)$ is 2-UDA iff at least n3n-3 states of $\rho_{A_{i}}$ are pure. It suffices to show the ”only if” part. Suppose only n4n-4 states of $\rho_{A_{i}}$ are pure, and, without loss of generality, they are $\rho_{A_{i}}=|\psi^{(i)}\rangle\!\langle\psi^{(i)}|_{A_{i}}$, for i=1,2,,n4i=1,2,...,n-4. Then any state $\sigma$ compatible with $\rho$ is $\sigma=\otimes_{i=1}^{n-4}|\psi^{(i)}\rangle\!\langle\psi^{(i)}|_{A_{i}}\otimes\sigma_{A_{n-3}...A_{n}}$. Then $\rho$ is 2-UDA iff $\sigma_{A_{n-3}...A_{n}}=(\rho_{A_{n-3}}\otimes\rho_{A_{n-2}})\otimes(D\oplus 0)$, where $\rho_{A_{n-3}}$ and $\rho_{A_{n-2}}$ are mixed. By Theorem 20, we obtain that $(\rho_{A_{n-3}}\otimes\rho_{A_{n-2}})\otimes(D\oplus 0)$ is not 2-UDA. Hence $\rho$ is not 2-UDA.
(iii) $\rho$ is 2-UDA only if $\rho_{A_{n-3}A_{n-2}}\otimes\rho_{A_{n-1}A_{n}}$ is 2-UDA. This derives three cases using Theorem 20. (a) $\rho_{A_{n-3}A_{n-2}}$ or $\rho_{A_{n-1}A_{n}}$ is pure, then $\rho$ is 2-UDA iff the rest part of the state is 2-UDA by (ii); (b) at least one of $\rho_{A_{j}}$ is pure, the form of $\rho$ reduces to (ii); (c) $\rho_{A_{n-3}A_{n-2}}$ and $\rho_{A_{n-1}A_{n}}$ are rank-two, $\rho_{A_{n-3}A_{n-2}}\otimes\rho_{A_{n-1}A_{n}}$ is LU equivalent to $(H\oplus 0)\otimes(D\oplus 0)$. By Theorem 22, $\rho_{A_{j}}\otimes\rho_{A_{n-3}A_{n-2}}\otimes\rho_{A_{n-1}A_{n}}$ is not 2-UDA if $\rho_{A_{j}}$ is rank-two, for j=1,2,,n4j=1,2,...,n-4. Hence, $\otimes_{j=1}^{n-4}\rho_{A_{j}}$ is 2-UDA, which implies that $\rho$ is 2-UDA iff $\rho_{A_{n-3}A_{n-2}}\otimes\rho_{A_{n-1}A_{n}}$ is 2-UDA. This completes the proof.   $\sqcap\sqcup$